### v2

# Estimating Causal Effects for Binary Outcomes Using Per-Decision Inverse Probability Weighting

Abstract
footnotetext: * To whom correspondence should be addressed.
Micro-randomized trials are commonly conducted for optimizing mobile health interventions such as push notifications for behavior change. In analyzing such trials, causal excursion effects are often of primary interest, and their estimation typically involves inverse probability weighting (IPW). However, in a micro-randomized trial additional treatments can often occur during the time window over which an outcome is defined, and this can greatly inflate the variance of the causal effect estimator because IPW would involve a product of numerous weights. To reduce variance and improve estimation efficiency, we propose a new estimator using a modified version of IPW, which we call “per-decision IPW”. It is applicable when the outcome is binary and can be expressed as the maximum of a series of sub-outcomes defined over sub-intervals of time. We establish the estimator’s consistency and asymptotic normality. Through simulation studies and real data applications, we demonstrate substantial efficiency improvement of the proposed estimator over existing estimators (relative efficiency up to 1.45 and sample size savings up to 31% in realistic settings). The new estimator can be used to improve the precision of primary and secondary analyses for micro-randomized trials with binary outcomes.
Keywords Causal excursion effect; Inverse probability weighting; Log relative risk; Micro-randomized trial; Per-decision EMEE; Per-decision importance weighting.
1 Introduction
Mobile health interventions that provides individual-level support has been undergoing fast growth [1]. Smartphone- or wearable-based messages and push notifications are used for promoting antiretroviral adherence, smoking cessation, and weight maintenance [2, 3, 4]. A novel experimental design, micro-randomized trial (MRT), was proposed in recent years for developing such interventions and has been increasingly used by domain scientists [5, 6, 7, 8]. “Micro-randomization” refers to the repeated randomization among treatment options (e.g., delivering a message or not) in the trial. Such randomization can occur 100s or 1000s of times within each participant. For example, in Sense2Stop, a smoking cessation MRT, randomization occurs every minute for 10 days. In HeartSteps II [9], a physical activity MRT, randomization occurs every 5 minutes for 90 days. We describe a typical MRT design in Section 2.1.
As a result, MRT data consist of longitudinal treatments and outcomes. These outcomes are called “proximal outcomes”, because they are measured in near-term after each randomization and are often the direct target of the interventions. The primary and secondary analyses of an MRT focus on the causal excursion effect of the treatment on the proximal outcome and the effect moderation by baseline variables and time-varying contexts. Causal excursion effects provide information on whether the intervention is worth further investigation (through assessing the overall effect) and, if so, how one might improve the intervention (through assessing effect moderation by contextual information). We formally introduce causal excursion effect in Section 2.2.
Because of the repeated randomization, multiple treatments can occur during the time window over which the proximal outcome is defined. This time window is referred to as the “proximal outcome window.” For example, in the Sense2Stop MRT, the proximal outcome is whether the individual experienced a stress episode during the 120 minutes following a randomization, whereas randomizations occur every minute. Thus, in the 120-minute proximal outcome window, there are 119 chances for additional treatments to occur. In the HeartSteps II MRT, the proximal outcome is whether the individual takes enough steps to constitute an activity bout during the 30 minutes following a randomization, whereas randomizations occur every 5 minutes. Existing estimators use inverse probability weighting (IPW) to account for the additional treatments in the proximal outcome window [10, 11]. These estimators effectively discard the data from the time points where additional treatments do occur during the proximal outcome window, and they upweight the data from the time points where additional treatments do not occur during the proximal outcome window. Such estimators can have high variance because (i) any time point where additional treatments occur in the proximal outcome window will not contribute to the estimation; and (ii) the product of the IP weights can be large for the time points where no additional treatments occur in the proximal outcome window. Consequently, a large sample may be necessary to identify promising interventions with statistical confidence.
Our main contribution is a more efficient estimator for such settings—MRTs where additional treatments may occur during the proximal outcome window. We focused on MRTs with binary proximal outcomes, which are common in practice [8, 4]. For such MRTs, the current state-of-the-art estimator was developed by [11]. We proposed a more efficient estimator for the causal excursion effect and established its consistency and asymptotic normality. The relative efficiency of the proposed estimator over that of [11] in most simulation scenarios ranges between 1.08 and 1.45, which roughly translates to savings in sample size between 7% and 31%.
Our estimator is applicable as long as the binary proximal outcome can be expressed as the maximum of a series of sub-outcomes defined over sub-intervals of time. This is often the case because a binary proximal outcome is often defined as whether a specific event occurred during a pre-specified time window. Examples of such events include a stress episode during a 120-minute window [3], a physical activity bout during a 30-minute window [12], or an interaction with a smartphone app during a 3-day window [13]. The key intuition for the efficiency improvement is to realize that the value of a binary proximal outcome is determined as soon as the event of interest occurs, and there is no need to wait till the end of the proximal outcome window to assign this value. Thus, in estimation one does not need to multiply together all the IP weights but only those weights before the occurrence of the event. Consequently, (i) some data points that would otherwise have been removed (due to being multiplied with a weight 0) are retained; and (ii) for the time points where no additional treatments occur in the proximal outcome window, the product of the IP weights becomes smaller (due to fewer terms in the product). This leads to the efficiency gain. We refer to this modification of IPW as “per-decision IPW”, because the inclusion of each weight is determined by whether the event has occurred by each decision point (i.e., a time point where randomization occurs). A related modification idea of the IP weights for efficiency improvement was proposed by [14] in the reinforcement learning literature for the setting with a single continuous outcome (i.e., the cumulative reward) that is the sum of sub-outcomes. Our method applies to a different setting where the outcome itself is longitudinal binary.
The rest of paper is organized as follows. In Section 1.1 we introduce the Drink Less MRT, our motivating example. In Section 2 we define the causal excursion effect using the potential outcomes framework and provide identifiability results. In Section 3 we present the new estimator and the asymptotic properties. In Section 4 we present results from simulation studies on consistency and efficiency of the proposed estimator. In Section 5 we illustrate the proposed method using the Drink Less MRT. Section 6 concludes with discussion.
1.1 Data Example: Drink Less MRT
Drink Less is a smartphone app to help users reduce harmful alcohol consumption [15, 16]. An MRT was conducted (embedded as one arm in a three-arm RCT) to assess and improve the effectiveness of a push notification intervention component that targets user engagement [13]. Each of the 349 participants in the MRT was randomized everyday at 8pm for 30 days, each time with a probability of 0.6 to receive an engagement intervention in the form of a push notification and a probability of 0.4 to receive nothing. The content of each push notification was randomly selected from a pool of messages that encourage the user to keep track of their drinking behavior using the app. The researchers have identified a large causal effect on app engagement within 1 hour of the intervention delivery and a smaller effect within 24 hours of the intervention delivery [17]. In this illustration of our method, we aim to explore whether the effect wears off within 72 hours of the intervention delivery. We define the binary proximal outcome as whether the user opens the app in the subsequent 72 hours after each randomization at 8pm. During the 72-hour proximal outcome window, there are two additional occurrences of randomization (at 8pm in the next two days), and the probability of at least one additional treatment being delivered in the proximal outcome window is 0.84 ($=1-0.4^{2}$). Therefore, on average 84% of the data points would be discarded if one were to use the estimator by [11]. Our proposed method will substantially improve the estimation efficiency.
2 Definitions and Assumptions
2.1 Notation and MRT Data Structure
Assume there are $n$ individuals in an MRT. Each individual is enrolled for $T$ time points at which treatments can be delivered. We will refer to these time points as decision points. In the presentation below we assume that $T$ is the same for all individuals, but our method generalizes to settings where $T$ varies across individuals. We use capital letters without subscript $i$ to denote observations of a generic individual. Denote the treatment assignment at decision point $t$ by $A_{t}$. For simplicity we assume that the treatment assignment is binary, with $A_{t}=1$ indicating treatment and 0 indicating no treatment. Denote observations collected after decision point $t-1$ and before decision point $t$ by vector $X_{t}$. Let $X_{T+1}$ denote the information observed after decision point $T$. As will be discussed shortly, the $X_{t}$’s include both covariates and outcomes. For each individual, the entire trajectory of observed information is $O=(X_{1},A_{1},X_{2},A_{2},\ldots,X_{T},A_{T},X_{T+1})$. We assume that data from different individuals are independently and identically distributed draws from an unknown distribution $P_{0}$. Unless noted otherwise, all expectations are taken with respect to $P_{0}$.
When designing the intervention (before the MRT starts), the researchers may deem it unsafe or unethical to deliver a push notification at certain decision points—for example, when the individual is driving. This is captured by the notion of “being (in)eligible to be randomized. (The same notion has been referred to as “being (un)available” in the literature [18, 19]) Formally, $X_{t}$ contains an indicator $I_{t}$, where $I_{t}=1$ if the individual is eligible to be randomized at decision point $t$ and $I_{t}=0$ otherwise. If $I_{t}$ = 0, $A_{t}=0$ deterministically.
We use overbar to denote a sequence of variables: for example, $\bar{A}_{t}=(A_{1},A_{2},\ldots,A_{t})$. Information accrued from an individual up to decision point $t$ is denoted by $H_{t}=(X_{1},A_{1},X_{2},A_{2},\ldots,$ $X_{t-1},A_{t-1},X_{t})=(\bar{X}_{t},\bar{A}_{t-1})$. At each $t$, $A_{t}$ is randomized with randomization probability depending on $H_{t}$, denoted by $p_{t}(H_{t})=P(A_{t}=1\mid H_{t})$.
We consider MRTs where the proximal outcome takes binary value and is defined over a window of length $\Delta$ (of decision points) following each decision point; the same proximal outcome was considered by [11]. Here $\Delta\geq 1$ is a fixed positive integer. This window is the “proximal outcome window” referred to in Section 1. Formally, suppose $Y_{t,\Delta}$ is a function of $H_{t+\Delta}$; i.e., the definition of the proximal outcome for decision point $t$ can depend on anything observed up to and including $X_{t+\Delta}$. For example, if $\Delta=1$, then $Y_{t,\Delta}$ is a function of $H_{t+1}$, i.e., is observed before $A_{t+1}$. If $\Delta=2$, then $Y_{t,\Delta}$ is a function of $H_{t+2}$, in which case additional treatments (precisely $A_{t+1}$ in this case) can occur before the value of $Y_{t,\Delta}$ is observed. Figure 1 depicts for the case of $\Delta=3$ a few consecutive decision points and the variables we defined, except for the $R$-variables which will be defined in Section 2.3.
Figure 1: Illustration of the data structure and the maximum property with $\Delta=3$.
In the Drink Less MRT, the decision point is everyday at 8pm. Each individual is in the study for $T=30$ days. $A_{t}$ is whether a push notification is delivered at 8pm on day $t$. $X_{t}$ include baseline variables and time-varying variables such as the number of days in the study and user’s previous interaction with the app. The researchers deemed that it is always appropriate to deliver push notifications, so $I_{t}=1$ for all $t$. The proximal outcome, $Y_{t,\Delta}$, is whether the individual logged in to the app in the subsequent 72 hours; thus $\Delta=3$.
2.2 Potential Outcomes and Causal Excursion Effects
For a mobile health intervention, causal quantities of interest include whether the intervention has an impact on an individual’s behavior and how this impact varies with contextual information such as the number of days in the study. To define the causal effects precisely we use the potential outcomes framework [20, 21]. Lower case letters represent instantiations (non-random values) of the corresponding capital letter random variable. For example, $a_{t}$ is an instantiation of treatment $A_{t}$. For every individual, denote by $X_{t}(\bar{a}_{t-1})$ the $X_{t}$ that would have been observed at decision point $t$ if the individual had been assigned a treatment sequence of $\bar{a}_{t-1}$ prior to $t$. The potential outcome of $H_{t}$ under $\bar{a}_{t-1}$ is
$H_{t}(\bar{a}_{t-1})=\{X_{1},a_{1},X_{2}(a_{1}),a_{2},X_{3}(\bar{a}_{2}),a_{3},\ldots,X_{t-1}(\bar{a}_{t-2}),a_{t-1},X_{t}(\bar{a}_{t-1})\}.$ (1)
The potential outcome of $Y_{t,\Delta}=y(H_{t+\Delta})$ is $Y_{t,\Delta}(\bar{a}_{t+\Delta-1})$.
We consider the causal excursion effect (CEE) of $A_{t}$ on $Y_{t,\Delta}$ on the log relative risk scale [11], defined as
$\text{CEE}\left\{S_{t}(\bar{A}_{t-1})\right\}=\log\frac{E\left\{Y_{t,\Delta}(\bar{A}_{t-1},1,\bar{0}_{\Delta-1})\mid S_{t}(\bar{A}_{t-1}),I_{t}(\bar{A}_{t-1})=1\right\}}{E\left\{Y_{t,\Delta}(\bar{A}_{t-1},0,\bar{0}_{\Delta-1})\mid S_{t}(\bar{A}_{t-1}),I_{t}(\bar{A}_{t-1})=1\right\}}.$ (2)
Here $\bar{0}_{\Delta-1}$ is a vector of 0’s of length $\Delta-1$; $(\bar{A}_{t-1},1,\bar{0}_{\Delta-1})$ and $(\bar{A}_{t-1},0,\bar{0}_{\Delta-1})$ are sequences of treatment assignments of length $t+\Delta-1$. $S_{t}(\bar{A}_{t-1})$ is a length-$p$ vector of summary variables from $H_{t}(\bar{A}_{t-1})$, usually chosen by the domain scientist based on the scientific question regarding effect modification. In Drink Less MRT, we will first set $S_{t}(\bar{A}_{t-1})=1$ to obtain the fully marginal effect (i.e., averaged over all participants and all days in the study) of a push notification on the app use in the subsequent 72-hour window. We will then set $S_{t}(\bar{A}_{t-1})=(1,\text{Day}_{t})$ to assess how the effect is moderated by days in the study.
The causal excursion effect (2) is commonly used in the primary and secondary analyses for MRTs [22, 4]. It differs from most causal effects under time-varying treatments such as marginal structural models and structural nested mean models [23, 24]. Instead of contrasting fixed treatment trajectories, (2) contrasts between two stochastic treatment trajectories, $(\bar{A}_{t-1},1,\bar{0}_{\Delta-1})$ and $(\bar{A}_{t-1},0,\bar{0}_{\Delta-1})$. That is, (2) is a causal contrast between two excursions from the treatment policy: following the treatment policy up to time $t-1$ and then deviating from the policy to assign $A_{t}=1$ and no treatment for the next $\Delta-1$ decision points, and following the treatment policy up to time $t-1$ and then deviating to assign $A_{t}=0$ and no treatment for the next $\Delta-1$ decision points. (2) is conditional on $I_{t}(\bar{A}_{t-1})=1$ and $S_{t}(\bar{A}_{t-1})$, meaning that it concerns the subpopulation who, after following the treatment policy up to time $t-1$, are eligible to be randomized at $t$ and within a stratum defined by $S_{t}(\bar{A}_{t-1})$. In practice we often impose a model on how $\text{CEE}\left\{S_{t}(\bar{A}_{t-1})\right\}$ depends on $S_{t}(\bar{A}_{t-1})$ to pool across $S_{t}$-strata.
$\text{CEE}\left\{S_{t}(\bar{A}_{t-1})\right\}$ depends on the treatment policy in the MRT, because the variables in $H_{t}(\bar{A}_{t-1})\setminus S_{t}(\bar{A}_{t-1})$ are marginalized over in the conditional expectations in (2). Such dependence is scientifically desirable for two reasons. First, the causal excursion effect approximates the treatment effect in a real-world implementation. A good MRT policy would already have incorporated implementation considerations such as burden and feasibility through the choice of randomization probability and availability criteria, so a feasible policy will not deviate too far from the MRT policy. Second, the causal excursion effect indicates the effective deviations from the MRT policy and how the policy might be improved. A fully marginal effect (by setting $S_{t}(\bar{A}_{t-1})=1$ in (2)) indicates whether a treatment is worth further investigation, and an effect modification analysis (by setting $S_{t}(\bar{A}_{t-1})$ to be certain time-varying covariates in (2)) indicates whether the MRT policy should be modified to depend on time-varying covariates. In addition, the dependence of $\text{CEE}\left\{S_{t}(\bar{A}_{t-1})\right\}$ on the MRT policy resembles the primary analysis in factorial designs and allows one to design trials with a higher power to detect meaningful effects. Related marginalization ideas were considered by [25, 26]. See [18], [11], and [19] for more discussion on causal excursion effect and its use in MRT. See [27] for a comprehensive comparison of various causal estimands under time-varying treatments.
2.3 Maximum Property of the Binary Proximal Outcome
We consider settings where the binary proximal outcome $Y_{t,\Delta}$ can be expressed as the maximum of a series of sub-outcomes defined over sub-intervals of time. Formally, let $R_{t+1}$ denote a sub-outcome—that is, a binary response observable after decision point $t$ and before $t+1$ and thus is a function of $X_{t+1}$. We assume that $Y_{t,\Delta}$ satisfies the maximum property:
$Y_{t,\Delta}=\max(R_{t+1},R_{t+2},\ldots,R_{t+\Delta}).$ (3)
That is, $Y_{t,\Delta}=1$ if and only if any of the sub-outcomes between $t$ and $t+\Delta$ is 1. Figure 1 shows how $Y_{t,\Delta}$ is related to the sub-outcomes when $\Delta=3$. We assume that $R_{t+\Delta}$ and $Y_{t,\Delta}$ are well-defined even when $t+\Delta>T+1$; this is satisfied when the researchers continue to follow up trial participants for $\Delta$ time points after decision point $T$.
Equation (3) holds if the binary outcome $Y_{t,\Delta}$ is whether an event of interest occurs between $t$ and $t+\Delta$ and such events are “instantaneous”. In the Drink Less MRT example with $\Delta=3$, the event is a user logging in to the app. Logging in between $t$ and $t+3$ is equivalent to logging in between $t$ and $t+1$, or between $t+1$ and $t+2$, or between $t+2$ and $t+3$. Thus, in this example (3) is satisfied with $R_{t+1}$ denoting whether the individual logging in to the app between 8pm of day $t$ and 7:59pm of day $t+1$, in which case we have $Y_{t,\Delta=3}=\max(R_{t+1},R_{t+2},R_{t+3})$.
If, however, $Y_{t,\Delta}$ is defined by whether an event is sustained for a period of time, then (3) may not hold. For example, in the HeartSteps II MRT, the binary proximal outcome is whether an activity bout occurred during the 30-minute window. Consider an activity bout to be a period of sustained activity of 90 seconds or longer. Suppose an individual started activity 45 seconds prior to decision point $t+1$, and concluded activity 45 seconds after $t+1$. In this case, their $Y_{t,\Delta=2}$ would be 1, but there is no straightforward definition of $R_{t+1}$ to satisfy (3). In fact, there is no activity bout of 90 seconds that is completely inside either of the sub-intervals $(t,t+1)$ and $(t+1,t+2)$. To generalize our method to such cases, we can introduce a double-subscript version of the sub-outcome $R$ along with a modified maximum property to determine the first decision point, within the proximal outcome window, at which the event has occurred. However, this generalization would make the notation too cumbersome without adding substantial novelty. Therefore, we focus on the current version of maximum property (3), and present how to modify the sub-outcomes and the maximum property to generalize our method in Supplementary Material A.
2.4 Identification of Parameters
To express the causal excursion effect using observed data, we assume the following.
Assumption 1 (Stable unit treatment value assumption (SUTVA))
The data observed are equal to the set of potential outcomes corresponding to the treatment assignment actually received, and one’s potential outcomes are not affected by others’ treatment assignments. That is, $X_{t}=X_{t}(\bar{A}_{t-1})$ for any $t\leq T$. This implies $R_{t}(\bar{A}_{t-1})=R_{t}$ and $Y_{t,\Delta}(\bar{A}_{t+\Delta-1})=Y_{t,\Delta}$.
Assumption 2 (Sequential ignorability)
For $1\leq t\leq T$, the potential outcomes $\{R_{j}(\bar{a}_{j-1}):\bar{a}_{j-1}\in\{0,1\}^{\otimes j-1},(t+1)\leq j\leq T\}$ are independent of $A_{t}$ given $H_{t},I_{t}=1$. This implies that $\{Y_{j,\Delta}(\bar{a}_{j+\Delta-1}):\bar{a}_{j+\Delta-1}\in\{0,1\}^{\otimes j+\Delta-1},t\leq j\leq T\}$ are independent of $A_{t}$ given $H_{t},I_{t}=1$.
Assumption 3 (Positivity)
$P(A_{t}=1\mid H_{t},I_{t}=1)$ is bounded away from 0 and 1.
Assumption 4 (Maximum property)
$Y_{t,\Delta}(\bar{a}_{t+\Delta-1})=\max(R_{t+1}(\bar{a}_{t}),R_{t+2}(\bar{a}_{t+1}),\ldots,R_{t+\Delta}(\bar{a}_{t+\Delta-1}))$ for $1\leq t\leq T$.
Assumption 1 is standard for causal inference settings without interference. It may be violated if the mobile health intervention incorporates community features or the participants are clustered, in which case a causal inference framework that allows interference should be used instead [28]. (There is no community feature in the Drink Less app, our motivating example.) Assumption 2 is always valid for MRTs because the randomization probability of $A_{t}$ depends (at most) on $H_{t}$ by study design. Assumption 3 is valid as long as the researchers do not set the randomization probability to exactly 0 or 1 when $I_{t}=1$. Assumption 4 is the potential outcomes version of the maximum property (3).
Let $\mathbbm{1}(\cdot)$ denote the indicator function, and define $10^{0}=$ 1. In the Supplementary Material C, we prove that under the above 4 assumptions the casual excursion effect in (2) can be expressed using observed data distribution:
$\text{CEE}\left\{S_{t}(\bar{A}_{t-1})\right\}=\log\frac{E\bigg{[}E\left\{\prod_{j=t+1}^{t+\Delta-1}\left[\frac{\mathbbm{1}(A_{j}=0)}{1-p_{j}\left(H_{j}\right)}\right]^{\mathbbm{1}(\max_{1\leq s\leq j-t}R_{t+s}=0)}Y_{t,\Delta}\mid A_{t}=1,H_{t},I_{t}=1\right\}\mid S_{t},I_{t}=1\bigg{]}}{E\bigg{[}E\left\{\prod_{j=t+1}^{t+\Delta-1}\left[\frac{\mathbbm{1}(A_{j}=0)}{1-p_{j}\left(H_{j}\right)}\right]^{\mathbbm{1}(\max_{1\leq s\leq j-t}R_{t+s}=0)}Y_{t,\Delta}\mid A_{t}=0,H_{t},I_{t}=1\right\}\mid S_{t},I_{t}=1\bigg{]}}.$ (4)
We define $\prod_{j=t+1}^{t+\Delta-1}\left[\frac{\mathbbm{1}(A_{j}=0)}{1-p_{j}\left(H_{j}\right)}\right]^{\mathbbm{1}(\max_{1\leq s\leq j-t}R_{t+s}=0)}=1$ if $\Delta=1$. Equation (4) will be used to establish the asymptotic property of our estimator. We will slightly abuse the notation and denote the right hand side of (4) by $\text{CEE}(S_{t})$.
3 An Estimator Using Per-Decision Inverse Probability Weights
Consider a linear model for the causal excursion effect in (4):
$\text{CEE}(S_{t})=S_{t}^{T}\beta\quad\text{{for all} }1\leq t\leq T,$ (5)
where $\beta$ is $p$-dimensional. Assume the first entry of $S_{t}$ is 1 so that $\beta$ contains the intercept. For example, when $S_{t}=1$, $\text{CEE}(S_{t})=\beta_{0}$ is the fully marginal effect. When $S_{t}=(1,Z_{t})$ for some time-varying covariate $Z_{t}$, $\text{CEE}(S_{t})=\beta_{0}+\beta_{1}Z_{t}$, where $\beta_{0}$ denotes the effect when $Z_{t}=0$ and $\beta_{1}$ denotes the increase in the effect when $Z_{t}$ increases by 1. Model (5) also accommodates flexible time-varying effects when $S_{t}$ contains basis functions. We assume that (5) is a correct model and our goal is to estimate the unknown $p$-dimensional vector $\beta$.
Our estimator for $\beta$, denoted by $\hat{\beta}$, requires a working model for $E\{Y_{t,\Delta}(\bar{A}_{t-1},0,\bar{0}_{\Delta-1})\mid H_{t},I_{t}=1\}$, which characterizes the expected outcome under no treatment between $t$ and $t+\Delta$ given the history $H_{t}$ and $I_{t}=1$. We use $\exp\{g(H_{t})^{T}\alpha\}$ as the working model, where $g(H_{t})$ is a vector of features constructed from $H_{t}$ and $\alpha$ is a $q$-dimensional nuisance parameter. As will be established in Theorem 1, $\hat{\beta}$ is consistent and asymptotically normal regardless of whether the working model $\exp\{g(H_{t})^{T}\alpha\}$ is a correct model for $E\{Y_{t,\Delta}(\bar{A}_{t-1},0,\bar{0}_{\Delta-1})\mid H_{t},I_{t}=1\}$.
Our estimator $\hat{\beta}$, along with the nuisance estimator $\hat{\alpha}$, is obtained by solving the estimating equation $\frac{1}{n}\sum_{i=1}^{n}U_{i}(\alpha,\beta)=0$, with the estimating function $U_{i}(\alpha,\beta)$ defined as
$U_{i}(\alpha,\beta)=\sum_{t=1}^{T}I_{it}e^{-A_{it}S_{it}^{T}\beta}\epsilon_{it}(\alpha,\beta)M_{it}W_{it}\begin{bmatrix}g(H_{it})\\ \{A_{it}-\tilde{p}_{t}(S_{it})\}S_{it}\end{bmatrix}.$ (6)
The estimating function is multiplied by $I_{it}$ because the causal excursion effect only concerns the decision points where individuals are eligible to be randomized. The factor $e^{-A_{it}S_{it}^{T}\beta}$ is analogous to the blipping-down technique in multiplicative structural nested mean models [23]. The error term $\epsilon_{it}(\alpha,\beta)$ is defined as $\epsilon_{it}(\alpha,\beta)=Y_{it,\Delta}-e^{g(H_{it})^{T}\alpha+A_{it}S_{it}^{T}\beta}$, where $Y_{it,\Delta}$ is the proximal outcome $Y_{t,\Delta}$ for the $i$-th individual. The $M_{it}$ is defined as $M_{it}=\left\{\frac{\tilde{p}_{t}(S_{it})}{p_{t}(H_{it})}\right\}^{A_{it}}\left\{\frac{1-\tilde{p}_{t}(S_{it})}{1-p_{t}(H_{it})}\right\}^{1-A_{it}}$, which is similar to the stabilized IPW [24] and is employed by other estimators for causal excursion effects [18, 11]. Here $\tilde{p}_{t}(S_{it})\in(0,1)$ is similar to the numerator of a stabilized weight and can be chosen to improve efficiency by making $\tilde{p}_{t}(S_{it})$ as close to $p_{t}(H_{it})$ as possible. For example, if $p_{t}(H_{it})$ is a constant or depends only on $S_{it}$, one can simply set $\tilde{p}_{t}(S_{it})=p_{t}(H_{it})$ which makes $M_{it}\equiv 1$; if $p_{t}(H_{it})$ depends on variables in $H_{it}\setminus S_{it}$, one should not set $\tilde{p}_{t}(S_{it})=p_{t}(H_{it})$ but can set $\tilde{p}_{t}(S_{it})$ to be the prediction from a logistic regression with $A_{it}$ being the response and $S_{it}$ being the predictor. The $W_{it}$ is defined as $W_{it}=\prod_{j=t+1}^{t+\Delta-1}\left\{\frac{\mathbbm{1}(A_{ij}=0)}{1-p_{j}(H_{ij})}\right\}^{\mathbbm{1}(\max_{1\leq s\leq j-t}R_{i,t+s}=0)}$ which is the per-decision inverse probability weight that we will elaborate on later.
We established the consistency and asymptotic normality of $\hat{\beta}$ in Theorem 1, the proof of which is in Supplementary Material D.
Theorem 1
Assume the causal effect model (5) and Assumptions 1-4 in Section 2.4 hold. Let $\beta^{*}$ denote the true value of $\beta$ corresponding to the data generating distribution $P_{0}$. Define $\dot{U}_{i}(\alpha,\beta)=\partial U_{i}(\alpha,\beta)/\partial(\alpha^{T},\beta^{T})$. Under regularity conditions, $\hat{\beta}$ is consistent for $\beta^{*}$ and $\sqrt{n}(\hat{\beta}-\beta^{*})$ is asymptotically normal with mean zero and variance-covariance matrix $\Sigma$, and $\Sigma$ can be consistently estimated by the lower block diagonal $(p\times p)$ entry of the matrix
$\left\{\frac{1}{n}\sum_{i=1}^{n}\dot{U}_{i}(\hat{\alpha},\hat{\beta})\right\}^{-1}\left\{\frac{1}{n}\sum_{i=1}^{n}U_{i}(\hat{\alpha},\hat{\beta})U_{i}(\hat{\alpha},\hat{\beta})^{T}\right\}\left\{\frac{1}{n}\sum_{i=1}^{n}\dot{U}_{i}(\hat{\alpha},\hat{\beta})\right\}^{-1,T}.$
For the estimand $\beta^{*}$, [11] proposed an estimator called the estimator for marginal excursion effect (EMEE). The EMEE is the same as our estimator except that $W_{it}$ is replaced by $W_{it}^{\prime}=\prod_{j=t+1}^{t+\Delta-1}\left\{\frac{\mathbbm{1}(A_{ij}=0)}{1-p_{j}\left(H_{ij}\right)}\right\}$. The $W_{it}^{\prime}$ is a product of $(\Delta-1)$ inverse probability weights (IPW) to account for the fact that the causal excursion effect (2) concerns excursions under treatment trajectory $A_{i,t+1}=\cdots=A_{i,t+\Delta-1}=0$. Let us examine how the EMEE weighs the $(i,t)$-decision point by $W_{it}^{\prime}$ and later contrast with our estimator. Case 1: suppose for some $t+1\leq j\leq t+\Delta-1$, $A_{ij}=1$ (Table 1, Examples 1a and 1b). Because this is different from the treatment trajectory for the excursion, the $(i,t)$-decision point is assigned weight $W_{it}^{\prime}=0$ by EMEE. Case 2: suppose $A_{i,t+1}=\cdots=A_{i,t+\Delta-1}=0$ (Table 1, Examples 2a and 2b). In this case $W_{it}^{\prime}>1$ and the $(i,t)$-decision point is upweighted to account for itself and similar decision points that fall into Case 1 and thus have weight 0.
The key innovation in our proposed estimator is $W_{it}$, which improves efficiency over EMEE. We call $W_{it}$ the per-decision inverse probability weight because the inclusion of each of the $(\Delta-1)$ IPW terms is determined by each decision point: for each $j$ from $t+1$ to $t+\Delta-1$, if $\max_{1\leq s\leq j-t}R_{i,t+s}=0$ then $\frac{\mathbbm{1}(A_{ij}=0)}{1-p_{j}\left(H_{ij}\right)}$ is included in $W_{it}$; otherwise it is not included. Therefore, $W_{it}$ includes in its product the IPW terms from $j=t+1$ to right before the first occurrence of some $R_{j^{*}}=1$. By not including the IPW terms after some $R_{j^{*}}=1$, the per-decision IPW does not account for the treatment assignments after $j^{*}$. This is appropriate because the maximum property (3) implies that the treatment assignments after $j^{*}$ doesn’t affect the value of $Y_{t,\Delta}$ (which has to equal 1). To see how our estimator improves efficiency over the EMEE, let us now examine how our estimator weighs the $(i,t)$-decision point by $W_{it}$. For Case 1 described above, if some $R_{j}=1$ occurs before the first $A_{k}=1$ (Table 1, Example 1b), the $(i,t)$-decision point will have a positive weight. For Case 2 described above, if the first $R_{j^{*}}=1$ occurs for $j^{*}<t+\Delta-1$ (Table 1, Example 2b), then $W_{it}$ will include fewer terms than $W_{it}^{\prime}$ and thus be smaller, because $W_{it}=\prod_{j=t+1}^{j^{*}}\frac{1}{1-p_{j}\left(H_{ij}\right)}<\prod_{j=t+1}^{t+\Delta-1}\frac{1}{1-p_{j}\left(H_{ij}\right)}=W_{it}^{\prime}$.
To summarize, by capitalizing on the maximum property (3), the per-decision IP weight $W_{it}$ makes use of some decision points that would be otherwise discarded because $W_{it}^{\prime}=0$ (Example 1b) and reduces the weight of some decision points (Example 2b), and this leads to the efficiency improvement. The efficiency improvement of our estimator is larger when a larger proportion of the decision points belong to the above two types, which happens when the proximal outcome window length $\Delta$ is large, when the randomization probability is large, or when the baseline success probability (i.e., the success probability of the proximal outcome under no treatment) is larger. When the proximal outcome window is shorter than the distance between two decision points so that $\Delta=1$, our estimator becomes equivalent to EMEE. In Supplementary Material E we provided a theoretical analysis for the relative efficiency between the proposed estimator and the EMEE under simplifying assumptions.
Table 1: Comparison between per-decision IP weight $W_{it}$ in our estimator and standard IP weight $W_{it}^{\prime}$ in EMEE in a few examples for the case of $\Delta=3$. Subscript $i$ for all variables are omitted.
$R_{t}$ $A_{t+1}$ $R_{t+1}$ $A_{t+2}$ $R_{t+2}$ $Y_{t,\Delta}$ $W_{t}$ $W_{t}^{\prime}$
Example 1a 0 0 0 1 0 0 0 0
Example 1b 0 0 1 1 0 1 $\frac{1}{1-p_{t+1}\left(H_{t+1}\right)}$ 0
Example 2a 0 0 0 0 0 0 $\frac{1}{1-p_{t+1}\left(H_{t+1}\right)}\cdot\frac{1}{1-p_{t+2}\left(H_{t+2}\right)}$ $\frac{1}{1-p_{t+1}\left(H_{t+1}\right)}\cdot\frac{1}{1-p_{t+2}\left(H_{t+2}\right)}$
Example 2b 0 0 1 0 0 1 $\frac{1}{1-p_{t+1}\left(H_{t+1}\right)}$ $\frac{1}{1-p_{t+1}\left(H_{t+1}\right)}\cdot\frac{1}{1-p_{t+2}\left(H_{t+2}\right)}$
Related techniques of reducing IP weights to improve efficiency were developed for other settings. In off-policy reinforcement learning, the per-decision importance sampling technique was proposed to construct IP weights related to ours by leveraging the fact that earlier outcomes cannot depend on later treatments [14]. In Cox marginal structural model, to determine a subject’s contribution to a risk-set at time $t$, a stabilized weight was proposed which involves IP weights not through the end of the follow-up but only up till time $t$ [24]. Those settings are analogous to a special case in our setting ($T=1$), i.e., when focusing on a single proximal outcome that decomposes into a sequence of sub-outcomes. Our contribution is to extend the per-decision IPW technique to estimating causal excursion effects on longitudinal proximal outcomes. In Supplementary Material B, we provide a detailed explication on how our method connects to and differs from the literature.
Additional remarks about the property of the proposed estimator are as follows.
Remark 1
[…]
Remark 2
(Small-sample correction for inference.) The sandwich estimator for the variance in Theorem 1 can be anti-conservative with small sample size. To correct this, we adopt two small-sample correction techniques: (i) Similar to the correction to the sandwich variance estimator by [29], we pre-multiply the vector of each individual’s residual, $\epsilon_{t}(\hat{\alpha},\hat{\beta})$, by the inverse of the identity matrix minus the leverage for that individual; (ii) Similar to the correction to the critical value for inference by [6], we use as the critical value the quantile of a $t$-distribution, $t^{-1}_{n-p-q}(1-\eta/2)$, where $\eta$ is the significance level and $p$ and $q$ are the dimensions of $\beta$ and $\alpha$, respectively.
4 Simulation
4.1 Overview
We studied the finite-sample performance of our proposed estimator and three competitors: the EMEE by [11], and two generalized estimating equations [30]. We will refer to our estimator as “pd-EMEE”, where “pd” stands for per-decision.
4.2 Generative Model
We set $I_{t}\equiv 1$ and $T=100$. The generative model depends on two parameters, $\Delta$ (length of the proximal outcome window) and $p_{a}$ (constant randomization probability). $A_{t}\in\{0,1\}$ was generated from $\text{Bernoulli}(p_{a})$. A scalar time-varying covariate, $Z_{t}\in\{0,1,2\}$, was generated independently from all previous variables (i.e., $Z_{t}\perp\{Z_{s},A_{s},R_{s+1}:1\leq s\leq t-1\}$) with
$P(Z_{t}=0)=0.5^{-1/2\Delta}/C,\quad P(Z_{t}=1)=1/C,\quad P(Z_{t}=2)=0.5^{1/2\Delta}/C,$ (7)
where $C=0.5^{-1/2\Delta}+0.5^{1/2\Delta}+1$. The sub-outcome $R_{t+1}$ given $A_{t}$ and $H_{t}$ was generated as
$P(R_{t+1}=0\mid A_{t}=0,H_{t})$ $=0.5^{(1.5-0.5Z_{t})/\Delta},$ (8)
$P(R_{t+1}=0\mid A_{t}=1,H_{t})$ $=\frac{1-\{1-0.5^{(1.5-0.5Z_{t})/\Delta}\cdot(3/C\cdot 0.5^{1/\Delta})^{\Delta-1}\}\cdot e^{0.1+0.2Z_{t}}}{(3/C\cdot 0.5^{1/\Delta})^{\Delta-1}}.$ (9)
We set $R_{t+1}=0$ for $T<t\leq T+\Delta$. The proximal outcome was generated as $Y_{t,\Delta}=\max(R_{t+1},\ldots,R_{t+\Delta})$.
In Supplementary Material F we showed that for $1\leq t\leq T-\Delta$,
$E(Y_{t,\Delta}(\bar{A}_{t-1},a_{t},\bar{0}_{\Delta-1})\mid H_{t}(\bar{A}_{t-1}),I_{t}(\bar{A}_{t-1})=1)=\{1-0.5^{(\Delta+0.5-0.5Z_{t})/\Delta}\cdot(3/C)^{\Delta-1}\}\cdot e^{a_{t}(0.1+0.2Z_{t})}.$ (10)
Therefore, $\text{CEE}(Z_{t})=0.1+0.2Z_{t}$ for $1\leq t\leq T-\Delta$ under the generative model, and this holds for all $\Delta\geq 1$ and $p_{a}\in(0,1)$. For $T-\Delta<t<T$, $\text{CEE}(Z_{t})$ is slightly different but its influence on the simulation results is negligible because in the simulations $\Delta\ll T$.
There is substantial variability in $R_{t+1}$ and $Y_{t,\Delta}$. When we vary $\Delta$ from 2 to 10 and vary $Z_{t}\in\{0,1,2\}$, (8) varies from 0.59 to 0.97, (9) varies from 0.38 to 0.80, and (10) varies from 0.41 to 0.58 when setting $a_{t}=0$ and from 0.57 to 0.80 when setting $a_{t}=1$.
We constructed this generative model based on three considerations. First, in order to fairly assess how the efficiency improvement is affected by $\Delta$ and $p_{a}$, we needed $\text{CEE}(Z_{t})$ to not depend on $\Delta$ or $p_{a}$. Thus, the distribution of the sub-outcome $R_{t}$ would depend on $\Delta$ because there are $\Delta$ sub-outcomes in $Y_{t,\Delta}$. Second, we wanted to construct a complicated $E(Y_{t,\Delta}(\bar{A}_{t-1},0,\bar{0}_{\Delta-1})\mid H_{t},I_{t}=1)$ to demonstrate the robustness of the estimator against misspecified working models. Third, we wanted $Z_{t}$ to depend on $\Delta$ as little as possible. In a real application, $Z_{t}$ (a covariate) would not depend on $\Delta$ (a choice in the analysis). However, we found it difficult to achieve while keeping the form of $\text{CEE}(Z_{t})$ simple. The current generating distribution of $Z_{t}$ (7) is the least dependent on $\Delta$ among the ones we came up with, and the three probabilities in (7) are all very close to 1/3 for $\Delta\geq 2$.
4.3 Simulation Result
We first conducted simulations with different $\Delta$: $\Delta=3$ and $\Delta=10$. We used a constant randomization probability $p_{a}=0.2$. We considered two estimands with different $S_{t}$: the first estimand is $\text{CEE}(S_{t})=\beta_{0}$ with $S_{t}=1$, and the second estimand is $\text{CEE}(S_{t})=\beta_{1}+\beta_{2}Z_{t}$ with $S_{t}=(1,Z_{t})$. They represent common primary and secondary analyses of MRTs: $\beta_{0}$ is the fully marginal causal excursion effect, and $\beta_{1}+\beta_{2}Z_{t}$ is the effect moderated by $Z_{t}$.
Four estimators were compared: our proposed pd-EMEE, the EMEE by [11], the log-linear generalized estimating equations with independent correlation structure (GEE.ind), and the log-linear generalized estimating equations with exchangeable correlation structure (GEE.exch). For pd-EMEE and EMEE, we used the correct model of CEE for both estimands and the incorrect working model $\exp\{g(H_{t})^{T}\alpha\}=\exp(\alpha_{0}+\alpha_{1}Z_{t})$ for $E(Y_{t,\Delta}(\bar{A}_{t-1},0,\bar{0}_{\Delta-1})\mid H_{t},I_{t}=1)$. For GEE.ind and GEE.exch, we use the marginal mean model $\exp(\alpha_{0}+\alpha_{1}Z_{t}+\beta_{0}A_{t})$ for the first estimand and $\exp\{\alpha_{0}+\alpha_{1}Z_{t}+A_{t}(\beta_{1}+\beta_{2}Z_{t})\}$ for the second estimand. The marginal mean models used by the GEEs are misspecified because the part involving the $\alpha$’s are incorrect. See Supplementary Material G for the exact form of the log-linear GEEs used.
When $\Delta=3$, the true parameter values are $\beta_{0}^{*}=0.283$ (obtained numerically by plugging (10) into (2)), $\beta_{1}^{*}=0.1$, and $\beta_{2}^{*}=0.2$. Both pd-EMEE and EMEE consistently estimate all parameters of interest with close to nominal confidence interval coverage (top halves of Tables 2 and 3). The confidence interval coverage is improved by the small sample correction (Remark 2). pd-EMEE is slightly more efficient than EMEE: for example, when $n=100$, the relative efficiency between pd-EMEE and EMEE for estimating $\beta_{0}$, $\beta_{1}$, and $\beta_{2}$ is $1.080.026^{2}/0.025^{2}=$ 1.08, $1.120.037^{2}/0.035^{2}=$ 1.12, and $1.150.029^{2}/0.027^{2}=$ 1.15, respectively. These values roughly translate to 7% ($=1-1/1.08$), 11%, and 13% savings in sample size if one were to power an MRT using pd-EMEE instead of EMEE. Both GEE.ind and GEE.exch are inconsistent because the marginal mean model is misspecified. The performance of $\hat{\alpha}_{0}$ and $\hat{\alpha}_{1}$ is not presented because $\alpha_{0}$ and $\alpha_{1}$ are nuisance parameters and thus not of interest.
When $\Delta=10$, the true parameter values are $\beta_{0}^{*}=0.304$, $\beta_{1}^{*}=0.1$, and $\beta_{2}^{*}=0.2$. The conclusions regarding consistency and confidence interval coverage are the same as when $\Delta=3$ (bottom halves of Tables 2 and 3). Here pd-EMEE is substantially more efficient than EMEE: for example, when $n=100$, the relative efficiency between pd-EMEE and EMEE for estimating $\beta_{0}$, $\beta_{1}$, and $\beta_{2}$ is $1.450.065^{2}/0.054^{2}=$ 1.45, $1.390.092^{2}/0.078^{2}=$ 1.39, and $1.400.064^{2}/0.054^{2}=$ 1.40, respectively, which roughly translates to 31%, 28%, and 29% savings in sample size. The efficiency improvement is greater compared to when $\Delta=3$.
Table 2: Performance of pd-EMEE, EMEE, GEE.ind and GEE.exch when estimating $\beta_{0}$ ($\Delta$ = 3 and 10). Based on 1,000 simulation replications.
Estimator $n$ Bias SD RMSE CP.unadj CP.adj
Setting: $\Delta=3$
30 0.006 0.045 0.045 0.94 0.96
50 0.004 0.037 0.037 0.93 0.94
pd-EMEE 100 0.005 0.025 0.026 0.94 0.94
30 0.005 0.047 0.048 0.95 0.96
50 0.004 0.039 0.039 0.93 0.95
EMEE 100 0.005 0.026 0.027 0.95 0.96
30 -0.061 0.032 0.069 0.51 0.54
50 -0.062 0.026 0.068 0.29 0.31
GEE.ind 100 -0.062 0.018 0.064 0.05 0.06
30 -0.063 0.032 0.071 0.48 0.52
50 -0.065 0.026 0.069 0.26 0.28
GEE.exch 100 -0.064 0.018 0.066 0.04 0.04
Setting: $\Delta=10$
30 0.026 0.103 0.106 0.94 0.96
50 0.022 0.084 0.087 0.94 0.95
pd-EMEE 100 0.022 0.054 0.058 0.95 0.96
30 0.032 0.127 0.130 0.94 0.96
50 0.026 0.099 0.103 0.94 0.96
EMEE 100 0.023 0.065 0.069 0.95 0.96
30 -0.168 0.025 0.169 0.00 0.00
50 -0.169 0.019 0.170 0.00 0.00
GEE.ind 100 -0.169 0.014 0.169 0.00 0.00
30 -0.175 0.024 0.177 0.00 0.00
50 -0.176 0.019 0.177 0.00 0.00
GEE.exch 100 -0.176 0.013 0.177 0.00 0.00
  • pd-EMEE, our proposed estimator; EMEE, the estimator in [11]; GEE.ind, generalized estimating equations with independent working correlation structure; GEE.exch, generalized estimating equations with exchangeable working correlation structure; SD, standard deviation; RMSE, root mean squared error; CP, $95\%$ confidence interval coverage probability before (unadj) and after (adj) small-sample correction.
Table 3: Performance of pd-EMEE, EMEE, GEE.ind and GEE.exch when estimating $(\beta_{1},\beta_{2})$ ($\Delta$ = 3 and 10). Based on 1,000 simulation replications.
$\beta_{1}$ $\beta_{2}$
Estimator Sample Size Bias SD RMSE CP(unadj) CP(adj) Bias SD RMSE CP(unadj) CP(adj)
Setting: $\Delta=3$
30 0.004 0.066 0.066 0.94 0.96 0.002 0.053 0.053 0.94 0.96
50 0.000 0.053 0.053 0.94 0.95 0.004 0.041 0.042 0.94 0.94
pd-EMEE 100 0.002 0.035 0.035 0.96 0.96 0.003 0.027 0.027 0.96 0.96
30 0.003 0.070 0.070 0.94 0.96 0.002 0.055 0.055 0.94 0.96
50 0.000 0.057 0.057 0.94 0.95 0.003 0.044 0.044 0.93 0.95
EMEE 100 0.002 0.037 0.037 0.96 0.97 0.003 0.029 0.029 0.95 0.96
30 -0.018 0.050 0.054 0.92 0.93 -0.038 0.040 0.056 0.81 0.84
50 -0.020 0.041 0.045 0.90 0.91 -0.039 0.032 0.050 0.74 0.75
GEE.ind 100 -0.018 0.026 0.032 0.91 0.91 -0.040 0.020 0.044 0.53 0.54
30 -0.019 0.050 0.054 0.92 0.93 -0.040 0.040 0.057 0.79 0.82
50 -0.020 0.041 0.045 0.90 0.91 -0.040 0.031 0.051 0.72 0.74
GEE.exch 100 -0.019 0.026 0.032 0.90 0.91 -0.041 0.020 0.046 0.50 0.51
Setting: $\Delta=10$
30 0.013 0.150 0.151 0.93 0.96 0.014 0.105 0.106 0.93 0.95
50 0.003 0.120 0.120 0.92 0.93 0.018 0.082 0.084 0.93 0.95
pd-EMEE 100 0.007 0.078 0.078 0.95 0.96 0.014 0.054 0.056 0.94 0.95
30 0.008 0.188 0.188 0.93 0.96 0.021 0.130 0.131 0.93 0.96
50 0.001 0.141 0.141 0.95 0.97 0.022 0.095 0.097 0.95 0.96
EMEE 100 0.006 0.092 0.092 0.97 0.97 0.015 0.064 0.066 0.96 0.96
30 -0.058 0.040 0.070 0.68 0.71 -0.109 0.029 0.112 0.05 0.06
50 -0.061 0.030 0.068 0.48 0.50 -0.107 0.022 0.110 0.00 0.00
GEE.ind 100 -0.060 0.021 0.063 0.21 0.22 -0.108 0.015 0.109 0.00 0.00
30 -0.061 0.039 0.073 0.63 0.68 -0.112 0.029 0.116 0.04 0.04
50 -0.063 0.030 0.070 0.42 0.44 -0.111 0.022 0.113 0.00 0.00
GEE.exch 100 -0.063 0.021 0.066 0.16 0.16 -0.112 0.015 0.113 0.00 0.00
  • pd-EMEE, our proposed estimator; EMEE, the estimator in [11]; GEE.ind, generalized estimating equations with independent working correlation structure; GEE.exch, generalized estimating equations with exchangeable working correlation structure. SD, standard deviation; RMSE, root mean squared error; CP, $95\%$ confidence interval coverage probability before (unadj) and after (adj) small-sample correction.
We then investigated how the relative efficiency between pd-EMEE and EMEE depends on $\Delta$ (proximal outcome window length) and $p_{a}$ (randomization probability) by varying $\Delta$ and $p_{a}$ separately. When $\Delta$ varies from 1 to 10 with $p_{a}=0.2$, the relative efficiency increases from 1 to 1.45 (left panel of Figure 2). When $p_{a}$ varies from 0.1 to 0.8 with $\Delta=3$, the relative efficiency increases from 1.05 to 1.76 (right panel of Figure 2). Therefore, a larger $\Delta$ or a larger $p_{a}$ makes the proposed pd-EMEE method even more efficient than the EMEE method. This phenomenon was expected and explained in Section 3.
Figure 2: Relative efficiency $\text{Var}(\hat{\beta}^{\text{EMEE}})/\text{Var}(\hat{\beta}^{\text{pd-EMEE}})$ against $\Delta$ (proximal outcome window length) and $p$ (randomization probability).
5 Application
We apply the proposed pd-EMEE to the Drink Less MRT data. Recall that there are $n=349$ participants in this MRT, each in the study for $T=30$ days. Each participant was randomized once every day at 8pm with 0.6 probability to receive a push notification encouraging using the Drink Less app ($A_{t}=1$) and with 0.4 probability to receive nothing ($A_{t}=0$). Individuals were eligible to be randomized at all decision points ($I_{t}\equiv 1$). The proximal outcome, $Y_{t,\Delta}$, is whether the individual logged in to the app in the subsequent 72 hours, so $\Delta=3$. The binary sub-outcome, $R_{t+1}$, is whether the individual logged in to the app between 8 pm of day $t$ and 7:59 pm of day $t+1$, and $Y_{t,\Delta}=\max(R_{t+1},R_{t+2},R_{t+3})$.
We used pd-EMEE and EMEE for two estimands: the fully marginal effect and the effect moderated by $\text{Day}_{t}$, the number of days in the study. $\text{Day}_{t}$ takes values $0,1,\ldots,29$. The fully marginal excursion effect, $\beta_{0}$, is defined by setting $S_{t}=1$ in (5):
$\beta_{0}=\log\frac{E\left\{Y_{t,3}(\bar{A}_{t-1},1,\bar{0}_{2})\right\}}{E\left\{Y_{t,3}(\bar{A}_{t-1},0,\bar{0}_{2})\right\}}.$
The effect moderated by $\text{Day}_{t}$, parameterized by $(\beta_{1},\beta_{2})$, is defined by setting $S_{t}=(1,\text{Day}_{t})$ in (5):
$\beta_{1}+\beta_{2}\text{Day}_{t}=\log\frac{E\left\{Y_{t,3}(\bar{A}_{t-1},1,\bar{0}_{2})\mid\text{Day}_{t}\right\}}{E\left\{Y_{t,3}(\bar{A}_{t-1},0,\bar{0}_{2})\mid\text{Day}_{t}\right\}}.$
$\beta_{1}$ is the causal excursion effect on the first day, and $\beta_{2}$ is the change in the effect with each additional day. The $g(H_{t})$ in the working model for $E\{Y_{t,3}(\bar{A}_{t-1},0,\bar{0}_{2})\mid H_{t},I_{t}=1\}$ include gender, age, employment type, the AUDIT score (a measure of alcohol misuse), and two time-varying variables about app usage: whether the individual used the app before 8 pm that day, and whether they used the app after 9 pm the previous day.
For each of $\beta_{0},\beta_{1}$, and $\beta_{2}$, pd-EMEE and EMEE yield estimates in the same direction, and pd-EMEE results in smaller standard errors and narrower confidence intervals (Table 4). For example, $\beta_{0}$ estimated using pd-EMEE is 0.127 with a 95% confidence interval (0.075, 0.180). This means that when averaged over the course of the 30-day study, a notification (followed by 2 days with no notifications) increases the chance of someone logging into the app in the next 72 hours by approximately 13.5% ($=e^{0.127}-1$). The same $\beta_{0}$ estimated using EMEE is 0.079 with a 95% confidence interval (-0.031, 0.188), about twice as wide as that obtained by pd-EMEE. The 95% confidence intervals for $\beta_{1}$ and $\beta_{2}$ obtained using pd-EMEE are also substantially narrower than those obtained using EMEE. This illustrates the substantial efficiency improvement by the proposed pd-EMEE over the EMEE.
Table 4: Estimated marginal excursion effect ($\beta_{0}$) and effect moderation by $\text{Day}_{t}$ (with model $\text{CEE}=\beta_{1}+\beta_{2}\text{Day}_{t}$) of Drink Less push notification on whether the user opens the app in the subsequent 72 hours.
Estimand Method Estimates SE 95% CI $p$-value
pd-EMEE 0.127 0.027 (0.075, 0.180) $<0.001$
$\beta_{0}$ : marginal excursion effect EMEE 0.079 0.056 (-0.031, 0.188) 0.159
pd-EMEE 0.195 0.046 (0.105, 0.286) $<0.001$
$\beta_{1}$: causal excursion effect on the first day EMEE 0.284 0.081 (0.123, 0.445) $<0.001$
pd-EMEE -0.006 0.003 (-0.012, 0.001) 0.092
$\beta_{2}$: change in the causal excursion effect with each additional day EMEE -0.016 0.005 (-0.026, -0.006) 0.002
  • SE, standard error; CI, confidence interval. SE, 95% CI and $p$-value are based on the small-sample correction described in Section 3, Remark 2.
6 Discussion
We proposed the per-decision estimator of marginal excursion effect (pd-EMEE) for the causal excursion effect of a time-varying treatment on a binary longitudinal outcome for micro-randomized trials. The proposed estimator is applicable as long as the binary outcome encodes the occurrence of an event of interest in a pre-specified time window. When this time window spans multiple decision points, the proposed estimator improves efficiency compared to the currently widely-used EMEE estimator [11], by using a modified form of inverse probability weighting (IPW) called per-decision IPW. This efficiency improvement can lead to more precise estimation of the causal effect or smaller sample size when planning an MRT. The efficiency improvement is greater when there is a higher chance of one or more additional treatments occurring in the proximal outcome window; for example, when the proximal outcome window consists of more decision points or when the randomization probability is larger. When the proximal outcome window doesn’t contain additional decision points (i.e., when the proximal outcome is measured immediately), the proposed method is equivalent to the EMEE in [11].
The method assumes that the binary proximal outcome satisfies a maximum property, i.e., the outcome can be expressed as the maximum of a series of sub-outcomes defined over sub-intervals of time. This holds when the outcome indicates the occurrence of an instantaneous event. For non-instantaneous events, if it is still possible to determine the first decision point, within the proximal outcome window, at which the binary outcome has already occurred, then one can use a similar estimator for efficiency improvement by assuming a generalized maximum property. The generalization is detailed in Supplementary Material A.
The estimator can also be generalized to settings where the number of decision points within a proximal outcome window is not fixed. Such a setting may arise if the decision points are irregularly spaced. In this setting, as long as the occurrence of the event may be determined early on in the proximal outcome window, one can still apply the per-decision IPW idea and thus discard the IPW terms after the event occurs to improve efficiency.
The maximum property may resemble the structure of a discrete-time survival outcome, and may raise the question of whether the proximal outcome should be treated as a survival outcome instead of a binary one. We believe the choice of the model and outcome type should depend on the scientific question. For some interventions, the goal is so that a desirable event, such as self-monitoring in Drink Less and activity bout in HeartStep, occurs within a reasonable time window but not necessarily as soon as possible, and thus the causal effect on the binary outcome contains is relevant for assessing and improving the intervention. For other settings, the goal might be to reduce the risk of a dangerous or harmful behavior as soon as possible, in which case the causal effect on the survival outcome might be more relevant. An example for the latter is problem anger prevention [31], where the intervention is to reduce the risk of aggressive behavior once anger is detected.
We considered the same causal excursion effect as in [11], where the future $\Delta-1$ treatment assignments ($A_{t+1},\ldots,A_{t+\Delta-1}$) are all set to 0. That is, we consider excursions starting from decision point $t$ that receives treatment (or no treatment) at $t$, and then receive no treatment for the next $\Delta-1$ decision points. As pointed out by a reviewer, the choice of this reference treatment regime depends on the scientific question, and in some applications other reference regimes may be of interest. In Supplementary Material H we provide an extension of the proposed method to a class of alternative reference regimes and evaluated the efficiency gain under those settings.
7 Software
R (R Development Core Team [32]) code to reproduce the simulations and the real data analysis can be downloaded at https://github.com/tqian/EMEE_binary_modified_weight. The Drink Less data set is publicly available at https://osf.io/mtcfa.
Acknowledgments
Lauren Bell is supported by a PhD studentship funded by the MRC Network of Hubs for Trials Methodology Research (MR/L004933/2-R18).
Supplementary Materials
Appendix A Generalized Version of Maximum Property
We present a generalized version of the maximum property and a double-subscript version of the variable $R$. Using this formulation, we can apply methodology developed in the paper to a broader range of problems as long as it is possible to determine the first decision point, within the proximal outcome window, at which the binary outcome has already occurred.
Define $R_{t,t+s}$ to be the indicator of whether the event of interest occurred between decision points $t$ and $t+s$. Under this definition, we have the following generalized maximum property:
$Y_{t,\Delta}=\max(R_{t,t+1},R_{t,t+2},\ldots,R_{t,t+\Delta}).$ (11)
The definition of $R_{t,t+s}$ and its relationship with $Y_{t,\Delta}$ is illustrated in Web Figure 3 for the case of $\Delta=3$. The estimator and the proofs work for this more general definition with minimal modification: One simply needs to replace $R_{t+s}$ by $R_{t,t+s}$ throughout.
For example, in the HeartSteps II MRT [12], the proximal outcome was whether there is an activity bout in the 30 minutes following a decision point. There was a decision point every 5 minutes, so we denote by $Y_{t,\Delta=6}$ the proximal outcome using our notation. Here $R_{t,t+1}$ is whether an activity bout occurred between $t$ and $t+1$, $R_{t,t+2}$ is whether an activity bout occurred between $t$ and $t+2$, and so on. Here (11) holds with $\Delta=6$.
For the rest of the supplementary material, we assume that the original maximum property (3) in the main paper holds.
Figure 3: Illustration of $R_{t,t+s}$ and its relationship with $Y_{t,\Delta}$ for the case of $\Delta=3$.
Appendix B Connection to the Literature
B.1 Connection between Per-Decision IPW (Our Approach) and the Per-Decision Importance Sampling
We describe a related IPW technique in the literature, the per-decision importance sampling (PDIS) in offline reinforcement learning [14], and discuss how our per-decision IPW approach connects to and differs from PDIS. To make connections to our approach, we describe a simple example of PDIS using a setting similar to MRT. See [33, Section 2] for the general form of PDIS. Suppose there are $\Delta$ decision points, and at each decision point a treatment is randomized with randomization probability $p_{t}$ and a short-term outcome $R_{t+1}$ is observed. Let a distal outcome (an outcome of interest) be $Y:=\sum_{t=1}^{\Delta}R_{t+1}$. Suppose the goal is to estimate $E\{Y(a,\bar{0}_{\Delta-1})\}$. Then we have
$E\{Y(a,\bar{0}_{\Delta-1})\}$ $=E\left[\left\{\frac{\mathbbm{1}(A_{1}=a)}{ap_{1}+(1-a)(1-p_{1})}\prod_{s=2}^{\Delta}\frac{\mathbbm{1}(A_{s}=0)}{1-p_{s}}\right\}\sum_{t=1}^{\Delta}R_{t+1}\right]$ (12)
$=E\left[\sum_{t=1}^{\Delta}\left\{\frac{\mathbbm{1}(A_{1}=a)}{ap_{1}+(1-a)(1-p_{1})}\prod_{s=2}^{t}\frac{\mathbbm{1}(A_{s}=0)}{1-p_{s}}\right\}R_{t+1}\right],$ (13)
where (12) is the identification result using the usual IPW and (13) is the identification result using PDIS. The intuition behind (13) is that earlier outcomes cannot depend on later treatments. The same intuition is used in our approach to derive the identification result for proximal outcomes satisfying the maximum property; i.e.,
$E\{Y_{t,\Delta}(\bar{A}_{t-1},a,\bar{0}_{\Delta-1})\mid S_{t}(\bar{A}_{t-1}),I_{t}(\bar{A}_{t-1})=1\}$
$=E\left(E\left\{\prod_{j=t+1}^{t+\Delta-1}\left[\frac{\mathbbm{1}(A_{j}=0)}{1-p_{j}\left(H_{j}\right)}\right]^{\mathbbm{1}(\max_{1\leq s\leq j-t}R_{t+s}=0)}Y_{t,\Delta}\mid A_{t}=a,H_{t},I_{t}=1\right\}\mid S_{t},I_{t}=1\right).$ (14)
The difference between our approach and PDIS is that (i) the PDIS literature typically considers the outcome to be a sum of sub-outcomes whereas we considers a binary outcome that is the maximum of sub-outcomes, (ii) the PDIS literature considers a single outcome at the end whereas we considered the setting where the main outcome itself is longitudinal (i.e., we consider $Y_{t,\Delta}$ for $1\leq t\leq T$), and (iii) the PDIS identification result (13) is solely based on IPW, whereas our identification result (14) is a combination of IPW and standardization.
B.2 Connection between Our Estimating Equation and the Literature
Our estimating function is defined as
$U_{i}(\alpha,\beta)=\sum_{t=1}^{T}I_{it}e^{-A_{it}S_{it}^{T}\beta}\epsilon_{it}(\alpha,\beta)M_{it}W_{it}\begin{bmatrix}g(H_{it})\\ \{A_{it}-\tilde{p}_{t}(S_{it})\}S_{it}\end{bmatrix}.$ (15)
The role of each term is discussed in Section 3. Per the request of a reviewer, here we explicate the derivation of our estimating function.
Besides the per-decision IP weight, $W_{it}$, which was explained in the previous subsection, our estimating function $U_{i}(\alpha,\beta)$ resembles several existing estimating functions in the literature: the multiplicative structural nested mean model (MSNMM) [23], the estimator for causal excursion effect (EMEE) [11], and the weighted and centered least squares (WCLS) estimator [18].
In particular, the WCLS estimating function for the case with continuous proximal outcome and $\Delta=1$ is
$\sum_{t=1}^{T}I_{it}\tilde{\epsilon}_{it}(\alpha,\beta)M_{it}\begin{bmatrix}g(H_{it})\\ \{A_{it}-\tilde{p}_{t}(S_{it})\}S_{it}\end{bmatrix},$
with $\tilde{\epsilon}_{it}(\alpha,\beta)=Y_{it,1}-g(H_{it}^{T}\alpha-(A_{it}-\tilde{p}_{t}(S_{it}))S_{it}^{T}\beta$. Our estimating function (15) adopts the weighting (by $M_{it}$) and the centering of the treatment (by $\tilde{p}_{t}(S_{it})$), which were used to (i) estimate a marginal causal effect instead of a fully conditional causal effect, and (ii) make the estimator robust to misspecified control variables (the $g_{t}(H_{it})$ part).
To extend the WCLS estimating function to binary outcome, the EMEE estimator [11] used a blipping-down technique of MSNMM developed by [23], i.e., pre-multiplying the residual term $\epsilon_{it}(\alpha,\beta)$ by a negative causal effect (to remove the causal effect, i.e., “blip” the outcome down to its counterpart under no treatment). We borrowed this idea from EMEE. This is where the term $e^{-A_{it}S_{it}^{T}\beta}$ comes from. Note that the original blipping-down technique [23] depends on an exponent representing the fully conditional causal effect (such as $e^{-A_{it}f(H_{it})^{T}\theta}$), but we used a blipping-down factor that represents the marginal causal effect because this is the causal effect of interest in our case.
Appendix C Proof of Identifiability Result
We use overbar and double subscript to denote the sequence of treatments between two time points: for example, for $t_{2}\geq t_{1}$, $\bar{A}_{t_{1}:t_{2}}:=(A_{t_{1}},A_{t_{1}+1},\ldots,A_{t_{2}})$. When there is no confusion, we will omit the subscript on $\bar{0}$: for example, $Y_{t,\Delta}(\bar{A}_{t-1},a,\bar{0})=Y_{t,\Delta}(\bar{A}_{t-1},a,\bar{0}_{\Delta-1})$.
Under Assumptions 1-4 in Section 2.4, we have
$E\left\{Y_{t,\Delta}(\bar{A}_{t-1},a,\bar{0})\mid A_{t}=a,H_{t},I_{t}=1\right\}$ (16)
$=$ $E\left\{\prod_{j=t+1}^{t+\Delta-1}\left[\frac{1(A_{j}=0)}{P(A_{j}=0\mid H_{j})}\right]^{1(\max_{1\leq s\leq j-t}R_{t+s}=0)}Y_{t,\Delta}\mid A_{t}=a,H_{t},I_{t}=1\right\},$ (17)
which implies the identification result (4) in the main paper. Equation (17) follows immediately from the statements of Lemma C.3 and Lemma C.4 below. We first define a set of new quantities used in the proof, then state and prove Lemmas C.1-C.4.
Consider any fixed $t$ with $1\leq t\leq m$. For each $11\leq u\leq\Delta-$ 1, let $C_{tu}(\bar{A}_{t-1},a,\bar{0}_{u-1})$ denote the event that the first nonzero $R_{t+s}(\bar{A}_{t-1},a,\bar{0})$ after time $t$ occurs at $s=u$; for notation simplicity we write $C_{tu}(\bar{A}_{t-1},a,\bar{0}_{u-1})$ as $C_{tu}$. That is, for $11\leq u\leq\Delta-$ 1
$C_{tu}:=\left\{\max_{1\leq s\leq u-1}R_{t+s}(\bar{A}_{t-1},a,\bar{0})=0\text{ and }R_{t+u}(\bar{A}_{t-1},a,\bar{0})=1\right\}.$
Let $C_{t\Delta}:=C_{t\Delta}(\bar{A}_{t-1},a,\bar{0}_{\Delta-1})$ denote the event $\left\{\max_{1\leq s\leq\Delta-1}R_{t+s}(\bar{A}_{t-1},a,\bar{0})=0\right\}$. (The definition of $C_{t\Delta}$ is different from the rest $C_{tu}$. This is intentional because, as we will see shortly, there is no need to distinguish between $R_{t+\Delta}(\bar{A}_{t-1},a,\bar{0})=1$ and $R_{t+\Delta}(\bar{A}_{t-1},a,\bar{0})=0$ for the purpose of proving identification. Also, omitting the potential treatment trajectory in $C_{tu}$ and $C_{t\Delta}$ will not cause confusion because we always condider the treatment trajectory $\bar{A}_{t-1},a,\bar{0}_{\Delta-1}$ in this section.)
Lemma C.1
If $\Delta\geq 2$, then for all $1\leq t\leq m$, for all $11\leq u\leq\Delta-$ 1, for all $a_{t}\in\{0,1\}$ and for all $\bar{a}_{t+u:t+\Delta-1}\in\{0,1\}^{\otimes(\Delta-u)}$,
$Y_{t,\Delta}(\bar{A}_{t-1},a_{t},\bar{0}_{\Delta-1})\times\mathbbm{1}(C_{tu})=Y_{t,\Delta}(\bar{A}_{t-1},a_{t},\bar{0}_{u-1},\bar{a}_{t+u:t+\Delta-1})\times\mathbbm{1}(C_{tu}).$ (18)
Proof 1 (Proof of Lemma C.1)
When $\mathbbm{1}(C_{tu})=0$, (18) holds trivially. When $\mathbbm{1}(C_{tu})=1$, it must be the case that $R_{t+u}(\bar{A}_{t-1},a_{t},\bar{0}_{u-1})=1$, and Assumption 4 implies that $Y_{t,\Delta}(\bar{A}_{t-1},a_{t},\bar{0}_{u-1},\bar{a}_{t+u:t+\Delta-1})$ and $Y_{t,\Delta}(\bar{A}_{t-1},a_{t},\bar{0}_{\Delta-1})$ both equal to 1. Thus (18) holds when the indicator function equals 1. This completes the proof.
Lemma C.2
Suppose Assumptions 1-4 hold. Given fixed $1\leq t\leq m$ and $1\leq u\leq\Delta$, for all $1\leq k\leq u$ we have
$E\left\{Y_{t,\Delta}(\bar{A}_{t-1},a,\bar{0})\mid A_{t}=a,H_{t},I_{t}=1,C_{tu}\right\}$
$=$ $E\left\{\prod_{j=t+1}^{t+k-1}\frac{1(A_{j}=0)}{P(A_{j}=0\mid H_{j})}\times Y_{t,\Delta}(\bar{A}_{t-1},a,\bar{0})\mid A_{t}=a,H_{t},I_{t}=1,C_{tu}\right\}.$
Note that for $k=1$, we define $\prod_{j=t+1}^{t+k-1}\frac{1(A_{j}=0)}{p(A_{j}=0\mid H_{j})}=1$.
Proof 2 (Proof of Lemma C.2)
We prove by induction. $k=1$ holds trivially. Now suppose $k=k_{0}$ for some $11\leq k_{0}\leq u-$ 1 holds and we are to show $k=k_{0}+1$ holds. We have
$E\left\{Y_{t,\Delta}(\bar{A}_{t-1},a,\bar{0})\mid A_{t}=a,H_{t},I_{t}=1,C_{tu}\right\}$
$=$ $E\left\{\prod_{j=t+1}^{t+k_{0}-1}\frac{\mathbbm{1}(A_{j}=0)}{p(A_{j}=0\mid H_{j})}\times Y_{t,\Delta}(\bar{A}_{t-1},a,\bar{0})\mid A_{t}=a,H_{t},I_{t}=1,C_{tu}\right\}$ (19)
$=$ $E\left[E\left\{\prod_{j=t+1}^{t+k_{0}-1}\frac{\mathbbm{1}(A_{j}=0)}{p(A_{j}=0\mid H_{j})}\times Y_{t,\Delta}(\bar{A}_{t-1},a,\bar{0})\mid A_{t}=a,H_{t+k_{0}},I_{t}=1,C_{tu}\right\}\mid A_{t}=a,H_{t},I_{t}=1,C_{tu}\right]$ (20)
$=$ $E\left[E\left\{\prod_{j=t+1}^{t+k_{0}-1}\frac{\mathbbm{1}(A_{j}=0)}{p(A_{j}=0\mid H_{j})}\times Y_{t,\Delta}(\bar{A}_{t-1},a,\bar{0})\mid A_{t}=a,H_{t+k_{0}},I_{t}=1,C_{tu}\right\}\right.$
×$\left.\times E\left\{\frac{\mathbbm{1}(A_{t+k_{0}}=0)}{P(A_{t+k_{0}}=0\mid A_{t}=a,H_{t+k_{0}},I_{t}=1,C_{tu})}\mid A_{t}=a,H_{t+k_{0}},I_{t}=1,C_{tu}\right\}\mid A_{t}=a,H_{t},I_{t}=1,C_{tu}\right]$ (21)
$=$ $E\left[E\left\{\prod_{j=t+1}^{t+k_{0}}\frac{\mathbbm{1}(A_{j}=0)}{p(A_{j}=0\mid H_{j})}\times Y_{t,\Delta}(\bar{A}_{t-1},a,\bar{0})\mid A_{t}=a,H_{t+k_{0}},I_{t}=1,C_{tu}\right\}\mid A_{t}=a,H_{t},I_{t}=1,C_{tu}\right]$ (22)
$=$ $E\left\{\prod_{j=t+1}^{t+k_{0}}\frac{\mathbbm{1}(A_{j}=0)}{p(A_{j}=0\mid H_{j})}\times Y_{t,\Delta}(\bar{A}_{t-1},a,\bar{0})\mid A_{t}=a,H_{t},I_{t}=1,C_{tu}\right\}.$ (23)
Here, (19) follows from the induction assumption, (20) and (23) follow from the law of iterated expectation, (21) follows from the fact that $E\left\{\frac{\mathbbm{1}(A_{t+k_{0}}=0)}{P(A_{t+k_{0}}=0\mid A_{t}=a,H_{t+k_{0}},I_{t}=1,C_{tu})}\mid A_{t}=a,H_{t+k_{0}},I_{t}=1,C_{tu}\right\}=1$ (valid under Assumptions 2 and 3), and (22) follows from Assumption 2 (in particular, conditional on $H_{t+k_{0}}$, $A_{t+k_{0}}$ is independent of all potential outcomes including $Y_{t,\Delta}$ and $C_{tu}$). This completes the proof.
Lemma C.3
We have
$E\left\{Y_{t,\Delta}(\bar{A}_{t-1},a,\bar{0})\mid A_{t}=a,H_{t},I_{t}=1\right\}$
$=$ $E\left\{\prod_{j=t+1}^{t+\Delta-1}\left[\frac{1(A_{j}=0)}{P(A_{j}=0\mid H_{j})}\right]^{1(\max_{1\leq s\leq j-t}R_{t+s}(\bar{A}_{t-1},a,\bar{0})=0)}Y_{t,\Delta}(\bar{A}_{t-1},a,\bar{0})\mid A_{t}=a,H_{t},I_{t}=1\right\}$ (24)
holds for all $t=1,\ldots,m$.
Proof 3 (Proof of Lemma C.3)
Because $\{C_{tu}:1\leq u\leq\Delta\}$ is a partition of the whole sample space, for any random variable $Z$ defined on the same sample space we have
$E\left\{Z\mid A_{t}=a,H_{t},I_{t}=1\right\}=\sum_{u=1}^{\Delta}E\left\{Z\mid A_{t}=a,H_{t},I_{t}=1,C_{tu}\right\}P\left\{C_{tu}\mid A_{t}=a,H_{t},I_{t}=1\right\}.$
Therefore, to prove (24) it suffices to show that for any $1\leq u\leq\Delta$,
$E\left\{Y_{t,\Delta}(\bar{A}_{t-1},a,\bar{0})\mid A_{t}=a,H_{t},I_{t}=1,C_{tu}\right\}$
$=$ $E\left\{\prod_{j=t+1}^{t+\Delta-1}\left[\frac{1(A_{j}=0)}{P(A_{j}=0\mid H_{j})}\right]^{1(\max_{1\leq s\leq j-t}R_{t+s}(\bar{A}_{t-1},a,\bar{0})=0)}Y_{t,\Delta}(\bar{A}_{t-1},a,\bar{0})\mid A_{t}=a,H_{t},I_{t}=1,C_{tu}\right\}.$ (25)
Conditional on $C_{tu}$ (for a fixed $1\leq u\leq\Delta$),
$\prod_{j=t+1}^{t+\Delta-1}\left[\frac{1(A_{j}=0)}{P(A_{j}=0\mid H_{j})}\right]^{1(\max_{1\leq s\leq j-t}R_{t+s}(\bar{A}_{t-1},a,\bar{0})=0)}=\prod_{j=t+1}^{t+u-1}\frac{1(A_{j}=0)}{P(A_{j}=0\mid H_{j})}.$
So right hand side of (25) equals
$E\left\{\prod_{j=t+1}^{t+u-1}\frac{1(A_{j}=0)}{P(A_{j}=0\mid H_{j})}\times Y_{t,\Delta}(\bar{A}_{t-1},a,\bar{0})\mid A_{t}=a,H_{t},I_{t}=1,C_{tu}\right\}.$ (26)
Display (26) equals $E\left\{Y_{t,\Delta}(\bar{A}_{t-1},a,\bar{0})\mid A_{t}=a,H_{t},I_{t}=1,C_{tu}\right\}$ by setting $k=u$ in Lemma C.2. This completes the proof.
Lemma C.4
Suppose Assumptions 1-4 hold. We have
$E\left\{\prod_{j=t+1}^{t+\Delta-1}\left[\frac{\mathbbm{1}(A_{j}=0)}{P(A_{j}=0\mid H_{j})}\right]^{\mathbbm{1}(\max_{1\leq s\leq j-t}R_{t+s}(\bar{A}_{t-1},a,\bar{0})=0)}Y_{t,\Delta}(\bar{A}_{t-1},a,\bar{0})\mid A_{t}=a,H_{t},I_{t}=1\right\}$
$=$ $E\left\{\prod_{j=t+1}^{t+\Delta-1}\left[\frac{\mathbbm{1}(A_{j}=0)}{P(A_{j}=0\mid H_{j})}\right]^{\mathbbm{1}(\max_{1\leq s\leq j-t}R_{t+s}=0)}Y_{t,\Delta}\mid A_{t}=a,H_{t},I_{t}=1\right\}.$ (27)
Proof 4 (Proof of Lemma C.4)
We prove Lemma C.4 by showing that (27) holds when conditioning both sides on $C_{tu}$ for each $1\leq u\leq\Delta$. Lemma C.4 then follows because $\{C_{tu}:1\leq u\leq\Delta\}$ forms a partition of the whole sample space.
The left hand side of (27) conditional on $C_{t\Delta}$ is
$E\left\{\prod_{j=t+1}^{t+\Delta-1}\left[\frac{\mathbbm{1}(A_{j}=0)}{P(A_{j}=0\mid H_{j})}\right]^{\mathbbm{1}(\max_{1\leq s\leq j-t}R_{t+s}(\bar{A}_{t-1},a,\bar{0})=0)}Y_{t,\Delta}(\bar{A}_{t-1},a,\bar{0})\mid A_{t}=a,H_{t},I_{t}=1,C_{t\Delta}\right\}$
$=$ $E\left\{\prod_{j=t+1}^{t+\Delta-1}\left[\frac{\mathbbm{1}(A_{j}=0)}{P(A_{j}=0\mid H_{j})}\right]Y_{t,\Delta}(\bar{A}_{t-1},a,\bar{0}_{\Delta-1})\mid A_{t}=a,H_{t},I_{t}=1,C_{t\Delta}\right\}$ (28)
$=$ $E\left\{\prod_{j=t+1}^{t+\Delta-1}\left[\frac{\mathbbm{1}(A_{j}=0)}{P(A_{j}=0\mid H_{j})}\right]Y_{t,\Delta}(\bar{A}_{t-1},a,\bar{A}_{t+1:t+\Delta-1})\mid A_{t}=a,H_{t},I_{t}=1,C_{t\Delta}\right\}$ (29)
$=$ $E\left\{\prod_{j=t+1}^{t+\Delta-1}\left[\frac{\mathbbm{1}(A_{j}=0)}{P(A_{j}=0\mid H_{j})}\right]Y_{t,\Delta}\mid A_{t}=a,H_{t},I_{t}=1,C_{t\Delta}\right\},$ (30)
which is the right hand side of (27) conditional on $C_{t\Delta}$. Here, (28) follows from the definition of $C_{t\Delta}$; (29) follows from $\prod_{j=t+1}^{t+\Delta-1}\mathbbm{1}(A_{j}=0)Y_{t,\Delta}(\bar{A}_{t-1},a,\bar{0}_{\Delta-1})=\prod_{j=t+1}^{t+\Delta-1}\mathbbm{1}(A_{j}=0)Y_{t,\Delta}(\bar{A}_{t-1},a,\bar{A}_{t+1:t+\Delta-1})$, a consequence of Assumption 1; and (30) follows from Assumption 1.
The left hand side of (27) conditional on $C_{tu}$ for some $11\leq u\leq\Delta-$ 1 is
$E\left\{\prod_{j=t+1}^{t+\Delta-1}\left[\frac{\mathbbm{1}(A_{j}=0)}{P(A_{j}=0\mid H_{j})}\right]^{\mathbbm{1}(\max_{1\leq s\leq j-t}R_{t+s}(\bar{A}_{t-1},a,\bar{0})=0)}Y_{t,\Delta}(\bar{A}_{t-1},a,\bar{0})\mid A_{t}=a,H_{t},I_{t}=1,C_{tu}\right\}$
$=$ $E\left\{\prod_{j=t+1}^{t+u-1}\left[\frac{\mathbbm{1}(A_{j}=0)}{P(A_{j}=0\mid H_{j})}\right]Y_{t,\Delta}(\bar{A}_{t-1},a,\bar{0})\mid A_{t}=a,H_{t},I_{t}=1,C_{tu}\right\}$ (31)
$=$ $E\left\{\prod_{j=t+1}^{t+u-1}\left[\frac{\mathbbm{1}(A_{j}=0)}{P(A_{j}=0\mid H_{j})}\right]Y_{t,\Delta}(\bar{A}_{t-1},a,\bar{0}_{u-1},\bar{A}_{t+u:t+\Delta-1})\mid A_{t}=a,H_{t},I_{t}=1,C_{tu}\right\}$ (32)
$=$ $E\left\{\prod_{j=t+1}^{t+u-1}\left[\frac{\mathbbm{1}(A_{j}=0)}{P(A_{j}=0\mid H_{j})}\right]Y_{t,\Delta}(\bar{A}_{t+\Delta-1})\mid A_{t}=a,H_{t},I_{t}=1,C_{tu}\right\}$ (33)
$=$ $E\left\{\prod_{j=t+1}^{t+u-1}\left[\frac{\mathbbm{1}(A_{j}=0)}{P(A_{j}=0\mid H_{j})}\right]Y_{t,\Delta}\mid A_{t}=a,H_{t},I_{t}=1,C_{tu}\right\}$ (34)
$=$ $E\left\{\prod_{j=t+1}^{t+\Delta-1}\left[\frac{\mathbbm{1}(A_{j}=0)}{P(A_{j}=0\mid H_{j})}\right]^{\mathbbm{1}(\max_{1\leq s\leq j-t}R_{t+s}=0)}Y_{t,\Delta}\mid A_{t}=a,H_{t},I_{t}=1,C_{tu}\right\},$ (35)
which is the right hand side of (27) conditional on $C_{tu}$. Here, (31) follows from the definition of $C_{tu}$; (32) follows from Lemma C.1; (33) follows from $\prod_{j=t+1}^{t+u-1}\mathbbm{1}(A_{j}=0)Y_{t,\Delta}(\bar{A}_{t-1},a,\bar{0}_{u-1},\bar{A}_{t+u:t+\Delta-1})=\prod_{j=t+1}^{t+u-1}\mathbbm{1}(A_{j}=0)Y_{t,\Delta}(\bar{A}_{t-1},a,\bar{A}_{t+1:t+\Delta-1})$, a consequence of Assumption 1; (34) follows from Assumption 1; and (35) follows from the definition of $C_{tu}$. This completes the proof.
Appendix D Proof of Theorem 1
To establish Theorem 1, we assume the following regularity conditions.
Assumption D.1
Suppose $(\alpha,\beta)\in\Theta$, where $\Theta$ is a compact subset of a Euclidean space. Suppose there exists unique $(\alpha^{\prime},\beta^{\prime})\in\Theta$ such that $E[U_{i}(\alpha^{\prime},\beta^{\prime})]=0$.
Assumption D.2
Suppose $S_{t}$, $\exp(S_{t})$, $g(H_{t})$ and $\exp\{g(H_{t})\}$ all have finite fourth moment.
Assumption D.3
Suppose $U_{i}(\alpha,\beta)$ and $U_{i}(\alpha,\beta)U_{i}(\alpha,\beta)^{T}$ are both bounded by integrable functions.
Lemma D.1
Suppose Equation (5) and Assumptions 1-4 in the main paper hold. $\beta^{*}$ is the value of $\beta$ corresponding to the data generating distribution, $P_{0}$. For an arbitrary $g(H_{t})^{T}$ and an arbitrary $\alpha$, we have
$E\left[I_{t}e^{-A_{t}S_{t}^{T}\beta^{*}}\left\{Y_{t,\Delta}-e^{g\left(H_{t}\right)^{T}\alpha+A_{t}S_{t}^{T}\beta^{*}}\right\}M_{t}W_{t}\left\{A_{t}-\tilde{p}_{t}\left(S_{t}\right)\right\}S_{t}\right]=0.$
Proof 5 (Proof of Lemma D.1)
We have
$E\left[I_{t}e^{-A_{t}S_{t}^{T}\beta^{*}}\left\{Y_{t,\Delta}-e^{g\left(H_{t}\right)^{T}\alpha+A_{t}S_{t}^{T}\beta^{*}}\right\}M_{t}W_{t}\left\{A_{t}-\tilde{p}_{t}\left(S_{t}\right)\right\}S_{t}\right]$
$=$ $E\left(E\left[I_{t}e^{-A_{t}S_{t}^{T}\beta^{*}}\left\{Y_{t,\Delta}-e^{g\left(H_{t}\right)^{T}\alpha+A_{t}S_{t}^{T}\beta^{*}}\right\}M_{t}W_{t}\left\{A_{t}-\tilde{p}_{t}\left(S_{t}\right)\right\}S_{t}\mid H_{t}\right]\right)$ (36)
$=$ $E\left(E\left[e^{-A_{t}S_{t}^{T}\beta^{*}}\left\{Y_{t,\Delta}-e^{g\left(H_{t}\right)^{T}\alpha+A_{t}S_{t}^{T}\beta^{*}}\right\}M_{t}W_{t}\left\{A_{t}-\tilde{p}_{t}\left(S_{t}\right)\right\}\mid H_{t},I_{t}=1\right]I_{t}S_{t}\right)$ (37)
$=$ $E\left(E\bigg{[}e^{-S_{t}^{T}\beta^{*}}\left\{Y_{t,\Delta}-e^{g\left(H_{t}\right)^{T}\alpha+S_{t}^{T}\beta^{*}}\right\}\left\{1-\tilde{p}_{t}\left(S_{t}\right)\right\}\prod_{j=t+1}^{t+\Delta-1}\big{[}\frac{1\left(A_{j}=0\right)}{1-p_{j}\left(H_{j}\right)}\big{]}^{\mathbbm{1}(\max\limits_{1\leq s\leq j-t}R_{t+s}=0)}\mid H_{t},I_{t}=1,A_{t}=1\bigg{]}I_{t}\tilde{p}_{t}\left(S_{t}\right)S_{t}\right)$
$-E\left(E\bigg{[}\big{\{}Y_{t,\Delta}-e^{g\left(H_{t}\right)^{T}\alpha}\big{\}}\tilde{p}_{t}\left(S_{t}\right)\prod_{j=t+1}^{t+\Delta-1}\big{[}\frac{1\left(A_{j}=0\right)}{1-p_{j}\left(H_{j}\right)}\big{]}^{\mathbbm{1}(\max\limits_{1\leq s\leq j-t}R_{t+s}=0)}\mid H_{t},I_{t}=1,A_{t}=0\bigg{]}I_{t}\left\{1-\tilde{p}_{t}\left(S_{t}\right)\right\}S_{t}\right)$ (38)
$=$ $E\left[\left\{e^{-S_{t}^{T}\beta^{*}}E\Bigg{(}\prod_{j=t+1}^{t+\Delta-1}\bigg{[}\frac{\mathbbm{1}\left(A_{j}=0\right)}{1-p_{j}\left(H_{j}\right)}\bigg{]}^{\mathbbm{1}(\max\limits_{1\leq s\leq j-t}R_{t+s}=0)}Y_{t,\Delta}\mid H_{t},I_{t}=1,A_{t}=1\Bigg{)}\right.\right.$
$\left.\left.-E\Bigg{(}\prod_{j=t+1}^{t+\Delta-1}\bigg{[}\frac{\mathbbm{1}\left(A_{j}=0\right)}{1-p_{j}\left(H_{j}\right)}\bigg{]}^{\mathbbm{1}(\max\limits_{1\leq s\leq j-t}R_{t+s}=0)}Y_{t,\Delta}\mid H_{t},I_{t}=1,A_{t}=0\Bigg{)}\right\}I_{t}\tilde{p}_{t}\left(S_{t}\right)\left\{1-\tilde{p}_{t}\left(S_{t}\right)\right\}S_{t}\right]$ (39)
$=$ $0$ (40)
Here, (36) follows from the law of iterated expectation; (37) and (38) follows from expanding the expectation on $I_{t}$ and $A_{t}$ respectively; (39) follows from collecting terms; and (40) follows from the identifiability result (4) and the linear model (5) in the main paper, and the fact that $\beta^{*}$ is the true parameter value. This completes the proof.
Proof 6 (Proof of Theorem 1.)
Assumption D.1 implies that $(\hat{\alpha},\hat{\beta})\stackrel{{\scriptstyle p}}{{\to}}\left(\alpha^{\prime},\beta^{\prime}\right)$ by Theorem $5.9$ of [34]. Assumption D.1 (uniqueness of $\beta^{\prime}$) and Lemma D.1 imply that $\beta^{*}=\beta^{\prime}$. Therefore, $(\hat{\alpha},\hat{\beta})\stackrel{{\scriptstyle p}}{{\to}}\left(\alpha^{\prime},\beta^{*}\right)$, and in particular $\hat{\beta}$ is consistent for $\beta^{*}$.
Because $U_{i}(\alpha,\beta)$ is continuously differentiable and hence Lipschitz continuous, Theorem $5.21$ of [34] implies that $\sqrt{n}\big{\{}(\hat{\alpha},\hat{\beta})-\left(\alpha^{\prime},\beta^{*}\right)\big{\}}$ is asymptotically normal with mean zero and covariance matrix $\Sigma=\big{\{}E[\dot{U}_{i}(\alpha^{\prime},\beta^{*})]\big{\}}^{-1}E\left[U_{i}(\alpha^{\prime},\beta^{*})U_{i}(\alpha^{\prime},\beta^{*})^{T}\right]\big{\{}E[\dot{U}_{i}(\alpha^{\prime},\beta^{*})]\big{\}}^{-1,T}$.
Lastly, we show that this covariance matrix can be consistently estimated by
$\left\{\frac{1}{n}\sum_{i=1}^{n}\dot{U}_{i}(\hat{\alpha},\hat{\beta})\right\}^{-1}\left\{\frac{1}{n}\sum_{i=1}^{n}U_{i}(\hat{\alpha},\hat{\beta})U_{i}(\hat{\alpha},\hat{\beta})^{T}\right\}\left\{\frac{1}{n}\sum_{i=1}^{n}\dot{U}_{i}(\hat{\alpha},\hat{\beta})\right\}^{-1,T}.$
This is achieved by proving
$\frac{1}{n}\sum_{i=1}^{n}\dot{U}_{i}(\hat{\alpha},\hat{\beta})\stackrel{{\scriptstyle p}}{{\to}}E[\dot{U}_{i}(\alpha^{\prime},\beta^{*})]$ (41)
and
$\frac{1}{n}\sum_{i=1}^{n}U_{i}(\hat{\alpha},\hat{\beta})U_{i}(\hat{\alpha},\hat{\beta})^{T}\stackrel{{\scriptstyle p}}{{\to}}E\left[U_{i}(\alpha^{\prime},\beta^{*})U_{i}(\alpha^{\prime},\beta^{*})^{T}\right],$ (42)
combined with Slutsky’s theorem.
[…]
$~{}~{}\mathbbm{P}_{n}\{f_{i}(\hat{\alpha},\hat{\beta})\}-\mathbbm{P}\{f_{i}(\alpha,\beta)\}$
$=[\mathbbm{P}_{n}\{f_{i}(\hat{\alpha},\hat{\beta})\}-\mathbbm{P}\{f_{i}(\hat{\alpha},\hat{\beta})\}]+[\mathbbm{P}\{f_{i}(\hat{\alpha},\hat{\beta})\}-\mathbbm{P}\{f_{i}(\alpha,\beta)\}].$ (43)
The first term in (43) can be controlled by Glivenko-Cantelli Theorem:
$|\mathbbm{P}_{n}\{f_{i}(\hat{\alpha},\hat{\beta})\}|-\mathbbm{P}\{f_{i}(\hat{\alpha},\hat{\beta})\}\leq\sup_{\alpha,\beta}|(\mathbbm{P}_{n}-\mathbbm{P})f_{i}(\alpha,\beta)|=o_{p}(1),$
because whichever of $\dot{U}_{i}(\alpha,\beta)$ or $U_{i}(\alpha,\beta)U_{i}(\alpha,\beta)^{T}$ that we set $f_{i}$ to, it takes value in a Glivenko-Cantelli class as $(\alpha,\beta)$ is finite-dimensional and takes value in a compact set (Assumption D.1). The second term in (43), $\mathbbm{P}\{f_{i}(\hat{\alpha},\hat{\beta})\}-\mathbbm{P}\{f_{i}(\alpha,\beta)\}$, is $o_{p}(1)$ by the dominated convergence theorem because of $(\hat{\alpha},\hat{\beta})\stackrel{{\scriptstyle p}}{{\to}}\left(\alpha^{\prime},\beta^{*}\right)$ and Assumption D.3. Therefore, (43) is $o_{p}(1)$ for when $f_{i}=\dot{U}_{i}$ and when $f_{i}=U_{i}U_{i}^{T}$. Thus (41) and (42) are proven. This completes the proof of Theorem 1.
Appendix E Drivers of Efficiency Gain
To understand the drivers of the efficiency gain, here we analytically derive the relative efficiency between the proposed method (we call it “pd-EMEE” where “pd” stands for per-decision) and the EMEE method [11] under a set of simplifying assumptions. In particular, we assume that individuals are always available ($I_{it}\equiv 1$), the randomization probability $p_{t}(H_{it})$ is a constant $p$, the moderator $S_{it}=1$ (i.e., interested in the fully marginal effect), and $\tilde{p}_{t}(S_{it})$ is chosen to be equal to $p$. Furthermore, we consider the case $T=1$ with $\Delta\geq 1$ so for each individual there is only one proximal outcome $Y$ but $\Delta$ treatment occasions and $\Delta$ sub-outcomes $R_{2},\ldots,R_{\Delta+1}$. We assume that all the sub-outcomes are independent with all the treatments, the $\Delta$ sub-outcomes for the same individual are independent with each other, and the sub-outcomes are identically distributed: $R_{t}$ follows Bernoulli($q$) distribution for some $q$.
We consider a simplified variation of both the pd-EMEE and the EMEE methods. In both estimators, we assume that the nuisance parameter $\alpha$ is not estimated and is set to 0. In other words, no control variables are used in either estimator. This helps to further simplify the theoretical derivation while preserving the essence of this efficiency comparison, because both estimators are consistent for any control variables.
Under these simplifying assumptions, the estimating function for pd-EMEE ($m_{1}$) and the estimating function for EMEE ($m_{2}$) for the fully marginal causal excursion effect $\beta$ are as follows:
$m_{1}(\beta)$ $=e^{-A_{1}\beta}Y\prod_{j=2}^{\Delta}\frac{\mathbbm{1}(A_{j}=0)}{1-p}^{\mathbbm{1}(\max_{2\leq s\leq j}R_{j}=0)}(A_{1}-p),$
$m_{2}(\beta)$ $=e^{-A_{1}\beta}Y\prod_{j=2}^{\Delta}\frac{\mathbbm{1}(A_{j}=0)}{1-p}(A_{1}-p),$
and $\prod_{j=2}^{\Delta}(\cdot):=1$ if $\Delta=1$.
We derive in Section E.1 that under these simplifying assumptions, the asymptotic relative efficiency between the two estimators is (AVar denoting asymptotic variance)
[…]
The relative efficiency (44) equals 1 if $\Delta=1$ and increases with $\Delta$. When $\Delta>1$, the relative efficiency increases with $p$ and $q$ (Figure 4). This confirms our intuition: When $p$ is larger, additional treatments are more likely to occur in the proximal outcome window; when $q$ is larger, sub-outcomes are more likely to be equal to 1 in the proximal outcome window; when $\Delta$ is larger, additional treatments are more likely to occur in the proximal outcome window and sub-outcomes are more likely to be equal to 1 in the proximal outcome window. In all these situations, pd-EMEE can achieve a greater reduction on the terms in the IP weight product, thus improving efficiency.
Figure 4: Relative efficiency (44) against $p$, $q$, and $\Delta$. In each of the first two plots, the three lines for $\Delta=1$ are all flat and overlap with each other.
E.1 Proof of Equation (44)
Suppose all the simplifying assumptions stated in Section E hold.
Step 1: We derive the asymptotic variance of $\hat{\beta}^{\text{pd-EMEE}}$ that solves the following estimating equation:
$m_{1}(\beta)$ $=e^{-A_{1}\beta}Y\prod_{j=2}^{\Delta}\left\{\frac{\mathbbm{1}(A_{j}=0)}{1-p}\right\}^{\mathbbm{1}(\max_{2\leq s\leq j}R_{j}=0)}(A_{1}-p).$
The asymptotic variance equals
$\text{AVar}(\hat{\beta}^{\text{pd-EMEE}})=\{E(\dot{m}_{1})\}^{-1}E(m_{1}m_{1}^{T})\{E(\dot{m}_{1})\}^{-1,T}.$
We derive each of the terms in the product.
Step 1.1: Deriving $E(\dot{m}_{1})$. We have
$\dot{m}_{1}(\beta)=-e^{-A_{1}\beta}Y\prod_{j=2}^{\Delta}\left\{\frac{\mathbbm{1}(A_{j}=0)}{1-p}\right\}^{\mathbbm{1}(\max_{2\leq s\leq j}R_{j}=0)}A_{1}(A_{1}-p).$
For $2\leq u\leq\Delta$, let $B_{u}$ denote the event $R_{2}=\cdots=R_{u}=0$ and $R_{u+1}=1$; let $B_{1}$ denote the event $R_{2}=1$; and let $B_{0}$ denote the event that $R_{2}=\cdots=R_{\Delta+1}=0$. Note that $\{B_{t}\}_{0\leq t\leq\Delta}$ forms a partition of the whole sample space, and that $Y=0$ if and only if $B_{0}$ holds. We have
$E(\dot{m}_{1})$ $=-\sum_{t=0}^{\Delta}E\left[e^{-A_{1}\beta}Y\prod_{j=2}^{\Delta}\left\{\frac{\mathbbm{1}(A_{j}=0)}{1-p}\right\}^{\mathbbm{1}(\max_{2\leq s\leq j}R_{j}=0)}A_{1}(A_{1}-p)\mid B_{t}\right]P(B_{t})$
$=-\sum_{t=1}^{\Delta}E\left[e^{-A_{1}\beta}\prod_{j=2}^{t}\left\{\frac{\mathbbm{1}(A_{j}=0)}{1-p}\right\}A_{1}(A_{1}-p)\right]P(B_{t})$
$=-\sum_{t=1}^{\Delta}E\{e^{-A_{1}\beta}A_{1}(A_{1}-p)\}P(B_{t})$
$=-E\{e^{-A_{1}\beta}A_{1}(A_{1}-p)\}\sum_{t=1}^{\Delta}P(B_{t})$
$=-E\{e^{-A_{1}\beta}A_{1}(A_{1}-p)\}\{1-P(B_{0})\}$
$=-E\{e^{-A_{1}\beta}A_{1}(A_{1}-p)\}\{1-(1-q)^{\Delta}\},$
where the first equality follows from the fact that $\{B_{t}\}_{0\leq t\leq\Delta}$ forms a partition of the whole sample space, the second equality follows from the fact that $Y=0$ if and only if $B_{0}$ holds, the third equality follows from the law of iterated expectation and the fact that $p$ is the randomization probability of $A_{j}$, and the last equality follows from the assumption of i.i.d. $R_{j}$.
Step 1.2: Deriving $E(m_{1}m_{1}^{T})$. Because $S_{it}=1$, both $\beta$ and $m_{1}$ are scalars and thus $E(m_{1}m_{1}^{T})=E(m_{1}^{2})$. We have
$E(m_{1}m_{1}^{T})$ $=E\left[e^{-2A_{1}\beta}Y\prod_{j=2}^{\Delta}\left\{\frac{\mathbbm{1}(A_{j}=0)}{(1-p)^{2}}\right\}^{\mathbbm{1}(\max_{2\leq s\leq j}R_{j}=0)}(A_{1}-p)^{2}\right]$
$=\sum_{t=0}^{\Delta}E\left[e^{-2A_{1}\beta}Y\prod_{j=2}^{\Delta}\left\{\frac{\mathbbm{1}(A_{j}=0)}{(1-p)^{2}}\right\}^{\mathbbm{1}(\max_{2\leq s\leq j}R_{j}=0)}(A_{1}-p)^{2}\mid B_{t}\right]P(B_{t})$
$=\sum_{t=1}^{\Delta}E\left[e^{-2A_{1}\beta}\prod_{j=2}^{t}\left\{\frac{\mathbbm{1}(A_{j}=0)}{(1-p)^{2}}\right\}(A_{1}-p)^{2}\right]P(B_{t})$
$=\sum_{t=1}^{\Delta}E\left[e^{-2A_{1}\beta}(A_{1}-p)^{2}\right]P(B_{t})/(1-p)^{t-1}$
$=\sum_{t=1}^{\Delta}E\{e^{-2A_{1}\beta}(A_{1}-p)^{2}\}\frac{(1-q)^{t-1}q}{(1-p)^{t-1}}$
=$=\begin{cases}E\{e^{-2A_{1}\beta}(A_{1}-p)^{2}\}q\frac{1-\left(\frac{1-q}{1-p}\right)^{\Delta}}{1-\left(\frac{1-q}{1-p}\right)}&\text{if }p\neq q\\ E\{e^{-2A_{1}\beta}(A_{1}-p)^{2}\}q\Delta&\text{if }p=q,\end{cases}$
where the second equality follows from the fact that $\{B_{t}\}_{0\leq t\leq\Delta}$ forms a partition of the whole sample space, the third equality follows from the fact that $Y=0$ if and only if $B_{0}$ holds, the fourth equality follows from the law of iterated expectation and the fact that $p$ is the randomization probability of $A_{j}$, and the fifth equality follows from the assumption of i.i.d. $R_{j}$.
Step 1.3: Putting together the above two steps and we have
$\text{AVar}(\hat{\beta}^{\text{pd-EMEE}})$ =$=\{E(\dot{m}_{1})\}^{-1}E(m_{1}m_{1}^{T})\{E(\dot{m}_{1})\}^{-1,T}$
$=(1-(1-q)^{\Delta})^{-2}\cdot[E\{-e^{-A_{1}\beta}A_{1}(A_{1}-p)\}]^{-2}\cdot E\{e^{-2A_{1}\beta}(A_{1}-p)^{2}\}\cdot L,$ (45)
where
$L=\begin{cases}q\frac{1-\left(\frac{1-q}{1-p}\right)^{\Delta}}{1-\left(\frac{1-q}{1-p}\right)}&\text{if }p\neq q\\ q\Delta&\text{if }p=q.\end{cases}$
Step 2: We derive the asymptotic variance of $\hat{\beta}^{\text{EMEE}}$ that solves the following estimating equation:
$m_{2}(\beta)$ $=e^{-A_{1}\beta}Y\prod_{j=2}^{\Delta}\left\{\frac{\mathbbm{1}(A_{j}=0)}{1-p}\right\}(A_{1}-p).$
The asymptotic variance equals
$\text{AVar}(\hat{\beta}^{\text{EMEE}})=\{E(\dot{m}_{2})\}^{-1}E(m_{2}m_{2}^{T})\{E(\dot{m}_{2})\}^{-1,T}.$
We derive each of the terms in the product.
Step 2.1: Deriving $E(\dot{m}_{2})$. We have
$\dot{m}_{2}(\beta)=-e^{-A_{1}\beta}Y\prod_{j=2}^{\Delta}\left\{\frac{\mathbbm{1}(A_{j}=0)}{1-p}\right\}A_{1}(A_{1}-p).$
We have
$E(\dot{m}_{2})$ $=-E(Y)E\left[-e^{-A_{1}\beta}\prod_{j=2}^{\Delta}\left\{\frac{\mathbbm{1}(A_{j}=0)}{1-p}\right\}A_{1}(A_{1}-p)\right]$
$=-E(Y)E\{-e^{-A_{1}\beta}A_{1}(A_{1}-p)\}$
$=-(1-(1-q)^{\Delta})E\{-e^{-A_{1}\beta}A_{1}(A_{1}-p)\},$
where the first equality follows from the assumption that all the sub-outcomes are independent with all the treatments and thus $Y$ is independent of all the $A_{j}$’s, the second equality follows from the law of iterated expectation and the fact that $p$ is the randomization probability of $A_{j}$, and the last equality follows from the assumption of i.i.d. $R_{j}$.
Step 2.2: Deriving $E(m_{2}m_{2}^{T})$. Because $S_{it}=1$, both $\beta$ and $m_{2}$ are scalars and thus $E(m_{2}m_{2}^{T})=E(m_{2}^{2})$. We have
$E(m_{2}m_{2}^{T})$ $=E\left[e^{-2A_{1}\beta}Y\prod_{j=2}^{\Delta}\left\{\frac{\mathbbm{1}(A_{j}=0)}{(1-p)^{2}}\right\}(A_{1}-p)^{2}\right]$
$=E(Y)E\left[e^{-2A_{1}\beta}\prod_{j=2}^{\Delta}\left\{\frac{\mathbbm{1}(A_{j}=0)}{(1-p)^{2}}\right\}(A_{1}-p)^{2}\right]$
$=E(Y)E\{e^{-2A_{1}\beta}(A_{1}-p)^{2}\}/(1-p)^{\Delta-1}$
$=(1-(1-q)^{\Delta})E\{e^{-2A_{1}\beta}(A_{1}-p)^{2}\}/(1-p)^{\Delta-1},$
where the second equality follows from the assumption that all the sub-outcomes are independent with all the treatments and thus $Y$ is independent of all the $A_{j}$’s, the third equality follows from the law of iterated expectation and the fact that $p$ is the randomization probability of $A_{j}$, and the last equality follows from the assumption of i.i.d. $R_{j}$.
Step 2.3: Putting together the above two steps and we have
$\text{AVar}(\hat{\beta}^{\text{EMEE}})$ =$=\{E(\dot{m}_{2})\}^{-1}E(m_{2}m_{2}^{T})\{E(\dot{m}_{2})\}^{-1,T}$
$=(1-(1-q)^{\Delta})^{-1}\cdot[E\{-e^{-A_{1}\beta}A_{1}(A_{1}-p)\}]^{-2}\cdot E\{e^{-2A_{1}\beta}(A_{1}-p)^{2}\}/(1-p)^{\Delta-1}.$ (46)
Finally, combining (45) and (46) yields (44). This completes the proof.
Appendix F Proof of Equation (10) in Simulation
We show that
$E(Y_{t,\Delta}(\bar{A}_{t-1},a,\bar{0}_{\Delta-1})\mid H_{t}(\bar{A}_{t-1}),I_{t}(\bar{A}_{t-1})=1)=\{1-\gamma^{(\Delta+0.5(1-Z_{t}))/\Delta}\cdot(3/C)^{\Delta-1}\}\cdot e^{a(0.1+0.2Z_{t})}.$
Proof 7
Define $\phi_{0}(Z_{t})$ and $\phi_{1}(Z_{t})$ as follows:
$\phi_{0}(Z_{t}):=\gamma^{(1.5-0.5Z_{t})/\Delta}=P(R_{t+1}=0\mid A_{t}=0,H_{t})$
and
$\phi_{1}(Z_{t}):=\frac{1-\{1-\gamma^{(1.5-0.5Z_{t})/\Delta}\cdot(3/C\cdot\gamma^{1/\Delta})^{\Delta-1}\}\cdot\exp\{0.1+0.2Z_{t}\}}{(3/C\cdot\gamma^{1/\Delta})^{\Delta-1}}=P(R_{t+1}=0\mid A_{t}=1,H_{t})$
where $C=\gamma^{-1/2\Delta}+\gamma^{1/2\Delta}+1$. (In the generative model, the feasible value of $\gamma$ is subject to the constraint that $\phi_{0}(Z_{t})$ and $\phi_{1}(Z_{t})$ are bounded by $(0,1)$.) Then we have the following equation for $E(Y_{t,\Delta}(\bar{A}_{t-1},1,\bar{0})\mid H_{t})$:
$E(Y_{t,\Delta}(\bar{A}_{t-1},1,\bar{0})\mid H_{t})=P(Y_{t,\Delta}(\bar{A}_{t-1},1,\bar{0})=1\mid H_{t})$
$=1-P(R_{t+1}(\bar{A}_{t-1},1,\bar{0})=0,R_{t+2}(\bar{A}_{t-1},1,\bar{0})=0,\ldots,R_{t+\Delta}(\bar{A}_{t-1},1,\bar{0})=0\mid H_{t})$ (47)
$=1-P(R_{t+1}(\bar{A}_{t-1},1,\bar{0})=0\mid H_{t})\cdot P(R_{t+2}(\bar{A}_{t-1},1,\bar{0})=0\mid H_{t})\cdots P(R_{t+\Delta}(\bar{A}_{t-1},1,\bar{0})=0\mid H_{t}),$ (48)
where (47) follows from Assumption 4, and (48) follows from the fact that $R_{t+1},\ldots,R_{t+\Delta}$ are independent of each other in the generative model. Similarly, for $E(Y_{t,\Delta}(\bar{A}_{t-1},0,\bar{0})$ we have
$E(Y_{t,\Delta}(\bar{A}_{t-1},0,\bar{0})\mid H_{t})$
$=1-P(R_{t+1}(\bar{A}_{t-1},0,\bar{0})=0\mid H_{t})\cdot P(R_{t+2}(\bar{A}_{t-1},0,\bar{0})=0\mid H_{t})\cdots P(R_{t+\Delta}(\bar{A}_{t-1},0,\bar{0})=0\mid H_{t}).$ (49)
For any $11\leq j\leq\Delta-$ 1 and for any $a\in\{0,1\}$ we have
$P\left(R_{t+j+1}\left(A_{t-1},a,\bar{0}\right)=0\mid H_{t}\right)$
(Assumption 2) $=P\left(R_{t+j+1}\left(A_{t-1},a,\bar{0}\right)=0\mid H_{t},A_{t}=a\right)$
(law of iterated expectation) $=E\left\{E\left(\mathbbm{1}\left(R_{t+j+1}\left(\bar{A}_{t-1},a,\overline{0}\right)=0\right)\mid H_{t+j}\right)\mid H_{t},A_{t}=a\right\}$
$=E\{E\left(\mathbbm{1}\left(R_{t+j+1}=0\right)\left|H_{t+j},\left.A_{t+j}=0\right)\right|H_{t},A_{t}=a\right\}$ (50)
(definition of $\phi_{0}$) $=E\{\phi_{0}(Z_{t+j})\mid H_{t},A_{t}=a\}$
($\phi_{0}(Z_{t+j})$ is independent of $H_{t},A_{t}$) $=E\{\phi_{0}(Z_{t+j})\}$ (51)
$=\gamma^{(1.5-0.5\cdot 0)/\Delta}\cdot\gamma^{-1/2\Delta}/C+\gamma^{(1.5-0.5\cdot 1)/\Delta}\cdot 1/C+\gamma^{(1.5-0.5\cdot 2)/\Delta}\cdot\gamma^{1/2\Delta}/C$
$=3/C\cdot\gamma^{1/\Delta},$ (52)
where (50) follows from Assumption 1 and the fact that $R_{t+j+1}$ only depends on $(Z_{t+j},A_{t+j})$ in the generative model. Plugging (52) into (48) and we have
$E(Y_{t,\Delta}(\bar{A}_{t-1},1,\bar{0})\mid H_{t})$
$=1-\phi_{1}(Z_{t})\cdot\{3/C\cdot\gamma^{1/\Delta}\}^{\Delta-1}$
=$=\{1-\gamma^{(1.5-0.5Z_{t})/\Delta}\cdot(3/C\cdot\gamma^{1/\Delta})^{\Delta-1}\}\cdot\exp\{0.1+0.2Z_{t}\}$
=$=\{1-\gamma^{(\Delta+0.5-0.5Z_{t})/\Delta}\cdot(3/C)^{\Delta-1}\}\cdot\exp\{0.1+0.2Z_{t}\}$ (53)
Plugging (52) into (49) and we have
$E(Y_{t,\Delta}(\bar{A}_{t-1},0,\bar{0})\mid H_{t})$
$=1-\phi_{0}(Z_{t})\cdot\{3/C\cdot 0.5^{1/\Delta}\}^{\Delta-1}$
$=1-\gamma^{(1.5-0.5Z_{t})/\Delta}\cdot\{3/C\cdot\gamma^{1/\Delta}\}^{\Delta-1}$
$=1-\gamma^{(\Delta+0.5-0.5Z_{t})/\Delta}\cdot(3/C)^{\Delta-1}$ (54)
Based on (53) and (54), we have:
$E(Y_{t,\Delta}(\bar{A}_{t-1},A_{t},\bar{0})\mid H_{t},A_{t})=\{1-\gamma^{(\Delta+0.5-0.5Z_{t})/\Delta}\cdot(3/C)^{\Delta-1}\}\cdot\exp\{A_{t}(0.1+0.2Z_{t})\}$
Appendix G Generalized Estimating Equations (GEE) Used in Simulation
In the simulations, we considered two estimators based on generalized estimating equations (GEE): GEE with independent correlation structure (GEE.ind), and GEE with exchangeable correlation structure (GEE.exch). Here for concreteness we write down the mean and the variance models for both GEEs, and write down the estimating equation for GEE.ind.
Because we are considering causal effects on the relative risk scale, the GEE assumes the following mean model using a log link: $E(Y_{t,\Delta}\mid H_{t},A_{t})=e^{g(H_{t})^{T}\alpha+A_{t}S_{t}^{T}\beta}=:\mu_{t}(\alpha,\beta)$. Because $Y_{t,\Delta}$ is a binary variable, its variance is associated with the mean model: $\text{var}(Y_{t,\Delta}\mid H_{t},A_{t})=\mu_{t}(\alpha,\beta)\{1-\mu_{t}(\alpha,\beta)\}$. The estimating equation can be constructed following the original GEE proposal in [30]. Below we write down the estimating equation for GEE.ind:
$\mathbbm{P}_{n}\sum_{t}^{T}\frac{\partial\mu_{t}(\alpha,\beta)}{\partial(\alpha,\beta)^{T}}\text{var}(Y_{t,\Delta}\mid H_{t},A_{t})^{-1}\{Y_{t,\Delta}-\mu_{t}(\alpha,\beta)\}=0,$
which, by the form of $\mu_{t}(\alpha,\beta)$ and $\text{var}(Y_{t,\Delta}\mid H_{t},A_{t})$, simplifies to
$\mathbbm{P}_{n}\sum_{t}\left\{1-e^{g(H_{t})^{T}\alpha+A_{t}S_{t}^{T}\beta}\right\}^{-1}\left\{Y_{t,\Delta}-e^{g(H_{t})^{T}\alpha+A_{t}S_{t}^{T}\beta}\right\}\begin{bmatrix}g(H_{t})\\ A_{t}S_{t}\end{bmatrix}=0.$
Here we used $\mathbbm{P}_{n}$ to denote the empirical average over all individuals. In the simulation, the particular mean model used is $E(Y_{t,\Delta}\mid H_{t},A_{t})=e^{\alpha_{0}+\alpha_{1}Z_{t}+\beta A_{t}}$.
Appendix H Generalization to Other Reference Regimes
In the main paper, we considered the causal excursion effect contrasting treatment trajectories $(\bar{A}_{t-1},1,\bar{0}_{\Delta-1})$ and $(\bar{A}_{t-1},0,\bar{0}_{\Delta-1})$, where the future reference regime ($A_{t+1},\ldots,A_{t+\Delta-1}$) in the two trajectories are all set to 0. That is, we consider excursions starting from decision point $t$ that receives treatment (or no treatment) at $t$, and then receive no treatment for the next $\Delta-1$ decision points.
As pointed out by a reviewer, the choice of the reference treatment regime depends on the scientific question, and in some applications other reference regimes may be of interest. For example, in [18] they considered causal excursion effect contrasting treatment trajectories $(\bar{A}_{t-1},1,\bar{A}_{t+1:t+\Delta-1})$ and $(\bar{A}_{t-1},1,\bar{A}_{t+1:t+\Delta-1})$, where $\bar{A}_{t+1:t+\Delta-1}:=(A_{t+1},A_{t+2},\ldots,A_{t+\Delta-1})$. That is, they considered excursions starting from decision point $t$ that receive treatment (or no treatment) at $t$, and then switch back to follow the MRT policy for the next $\Delta-1$ decision points.
We generalize our method to reference regimes that are in between the reference regime we considered in the main paper and the reference regime considered in [18]. That is, excursions starting from decision point $t$ that receive treatment (or no treatment) at $t$, then receive no treatment for the next $K$ decision points (with some $10\leq K\leq\Delta-$ 1), then follow the MRT policy for the next $\Delta-1-K$ decision points.
Formally, consider the following causal excursion effect for $10\leq K\leq\Delta-$ 1:
$\text{CEE}^{(K)}\left\{S_{t}(\bar{A}_{t-1})\right\}=\log\frac{E\left\{Y_{t,\Delta}(\bar{A}_{t-1},1,\bar{0}_{K},\bar{A}_{t+K+1:t+\Delta-1})\mid S_{t}(\bar{A}_{t-1}),I_{t}(\bar{A}_{t-1})=1\right\}}{E\left\{Y_{t,\Delta}(\bar{A}_{t-1},0,\bar{0}_{K},\bar{A}_{t+K+1:t+\Delta-1})\mid S_{t}(\bar{A}_{t-1}),I_{t}(\bar{A}_{t-1})=1\right\}}.$ (55)
Under the same set of causal assumptions as in Section 2.4, we have the identification result:
$\text{CEE}^{(K)}\left\{S_{t}(\bar{A}_{t-1})\right\}=\log\frac{E\bigg{[}E\left\{\prod_{j=t+1}^{t+K}\left[\frac{\mathbbm{1}(A_{j}=0)}{1-p_{j}\left(H_{j}\right)}\right]^{\mathbbm{1}(\max_{1\leq s\leq j-t}R_{t+s}=0)}Y_{t,\Delta}\mid A_{t}=1,H_{t},I_{t}=1\right\}\mid S_{t},I_{t}=1\bigg{]}}{E\bigg{[}E\left\{\prod_{j=t+1}^{t+K}\left[\frac{\mathbbm{1}(A_{j}=0)}{1-p_{j}\left(H_{j}\right)}\right]^{\mathbbm{1}(\max_{1\leq s\leq j-t}R_{t+s}=0)}Y_{t,\Delta}\mid A_{t}=0,H_{t},I_{t}=1\right\}\mid S_{t},I_{t}=1\bigg{]}}.$ (56)
The difference between the two identification results (56) and (4) is that in (56) the product of weights is only up to decision point $t+K$ instead of up to $t+\Delta-1$, because after decision point $t+K$ the treatment trajectory comes back to the MRT protocol and thus no additional weight is needed.
Suppose we impose a linear model analogous to (5): $\text{CEE}^{(K)}(S_{t})=S_{t}^{T}\beta^{(K)}$. To estimate $\beta^{(K)}$, an estimating function analogous to (6) can be used by revising the weights:
$U_{i}^{(K)}(\alpha,\beta^{(K)})=\sum_{t=1}^{T}I_{it}e^{-A_{it}S_{it}^{T}\beta^{(K)}}\epsilon_{it}(\alpha,\beta^{(K)})M_{it}W_{it}^{(K)}\begin{bmatrix}g(H_{it})\\ \{A_{it}-\tilde{p}_{t}(S_{it})\}S_{it}\end{bmatrix},$ (57)
with
$W_{it}^{(K)}=\prod_{j=t+1}^{t+K}\left\{\frac{\mathbbm{1}(A_{ij}=0)}{1-p_{j}(H_{ij})}\right\}^{\mathbbm{1}(\max_{1\leq s\leq j-t}R_{i,t+s}=0)}.$
Using the same proof for Theorem 1, it is straightforward to show that consistency and asymptotic normality for $\hat{\beta}^{(K)}$ holds with the revised estimating function.
Below we present simulation results to compare this generalized version of our method (generalized pd-EMEE) and a generalized version of EMEE by [11], and see how the relative efficiency depends on the choice of reference regime. Here the EMEE by [11] needs to be generalized to handle the reference regimes we consider here. In particular, for the generalized version of EMEE, the estimating function would be the same as 57 except that $W_{it}^{(K)}$ would be replaced by
$\tilde{W}_{it}^{(K)}=\prod_{j=t+1}^{t+K}\left\{\frac{\mathbbm{1}(A_{ij}=0)}{1-p_{j}(H_{ij})}\right\}.$
We used the same generative model as in Section 4 and set $\Delta=10$. When $K$ varies from 0 to 9, the relative efficiency between the generalized pd-EMEE and the generalized EMEE increases from 1 to over 1.4 (Figure 5). This confirms the intuition that the farther the reference regime is from the MRT policy, the more efficiency gain there is when we use pd-EMEE over EMEE.
Figure 5: Relative efficiency between the generalized pd-EMEE and the generalized EMEE with various $K$, where $K$ parameterizes the reference regime. A relative efficiency of greater 1 means that the generalized pd-EMEE is more efficient.